### v2

# Quantum-limited superresolution of two arbitrary incoherent point sources: beating the resurgence of Rayleigh’s curse

Abstract
Superresolution has been demonstrated to overcome the limitation of the Rayleigh’s criterion and achieve significant improvement of the precision in resolving the separation of two incoherent optical point sources. However, in recent years, it was found that if the photon numbers of the two incoherent optical sources are unknown, the precision of superresolution vanishes when the two photon numbers are actually different. In this work, we first analyze the estimation precision of the separation between two incoherent optical sources with the same point-spread functions in detail, and show that when the photon numbers of the two optical sources are different but sufficiently close, the superresolution can still realized but with different precisions. We find the condition on how close the photon numbers of two optical sources need to be to realize the superresolution, and derive the precision of superresolution in different regimes of the photon number difference. We further consider the superresolution for two incoherent optical sources with different point-spread functions, and show that the competition between the difference of photon numbers, the difference of the two point-spread functions and the separation of the two optical sources determines the precision of superresolution. The results exhibit precision limits distinct from the case of two point sources with identical point-spread functions and equal photon numbers, and extend the realizable regimes of the quantum superresolution technique. The results are finally illustrated by Gaussian point-spread functions.
I Introduction
Optical imaging is an important sensing technique that has contributed to various research areas, such as medical technology, astronomy, nanotechnology and biochemistry, etc. Typical imaging methods include classical and quantum imaging [1, 2], fluorescence imaging [3, 4, 5], the superlens [6, 7], two-photon microscopy [8, 9], quantum holography [10, 11] and quantum lithography[12, 13], etc. However, by the Rayleigh’s criterion[14], when the separation of the two sources is smaller than the widths of two closely located point sources, it becomes extremely difficult to distinguish between the two sources. This imposes a severe limitation on the resolution of imaging.
The Rayleigh’s criterion indicates the insufficiency of direct measurement for high-resolution imaging in the positional space of optical sources. To precisely measure the separation of two point sources, it is critical to find a suitable measurement basis beyond the positional basis and a proper estimation method so that the imaging precision can surpass the classical precision limit of direct imaging. The quantum parameter estimation theory provides the mathematical tool to solve this problem. The quantum parameter estimation theory is rooted in its classical counterpart. The classical parameter estimation theory optimizes the estimation precision of unknown parameters over all possible estimation strategies and gives the optimal estimation precision for parameters in a given parameter-dependent probability distribution by the well-known Cramér-Rao’s bound. The quantum parameter estimation theory considers the parameters in a quantum system, and further optimizes the estimation precision over all possible measurement basis in addition to the optimization of estimation strategies, since different quantum measurement bases lead to different probability distributions of the measurement results and thus to different estimation precisions of the unknown parameters. The precision limit of quantum estimation is characterized by the quantum Cramér-Rao bound, which is determined by the inverse of the quantum Fisher information matrix [15, 16, 17, 18, 19].
Inspired by the quantum estimation theory, Tsang et al. proposed a technique of superresolution for two incoherent optical sources [20] to surpass the Rayleigh’s criterion. While the measurement precision of the separation vanishes by the classical direct imaging method when the separation is close to zero, which is known as Rayleigh’s curse, the precision of estimating the separation of the two sources does not vanish by optimizing the measurement basis. This is in sharp contrast to the Rayleigh’s criterion, which shows the power of quantum estimation theory and opens up a new avenue for precision resolution of two incoherent optical point sources.
The field of superresolution for incoherent optical sources has undergone a rapid and extensive development in recent years since Tsang et al.’s seminal work, in both theories and experimental implementations [21]. On one hand, the theory of multi-parameter estimation, essential to the task of resolving incoherent optical sources, has made significant progress in recent years, which has continuously stimulated the researches on the superresolution imaging [22, 23, 24, 25]. The superresolution theory has been extended for two-dimensional optical sources [26, 27, 28, 29], three-dimensional optical sources [30, 31], multiple point sources [32, 22, 33], thermal sources [34, 35, 36], far field and near field [37, 38, 39], dark field [40], and various types of optical media [41, 42, 43]. The influence of noises from the environments on the superresolution imaging has also been investigated [44, 45, 46, 47, 48, 49]. Moreover, the superresolution theory has been introduced to localizing and identifying objects [50, 51, 52, 53], discrete quantum imaging [54], spectral superresolution [55, 56], distributed quantum sensing [57, 58], superresolution microscopy [59, 60], etc.
In addition to the theoretical development, physical implementations of quantum superresolution have been proposed [61, 62, 63, 64, 65, 66, 67, 68] and experimentally realized in various ways, e.g., by optical interferometry [69, 70], with the assistance of phase information [71] or point-spread function engineering [72], using heterodyne detection [73], holographic technique [74], parity sorting [75], mode-selective measurement [76], spectrotemporal shaping [77], etc. Meanwhile, quantum superresolution has also been applied to optical localization and microscopy [78, 79, 80, 81, 82], imaging of remote bodies and stars [83, 84], etc.
An implicit assumption in the original superresolution protocol is that the intensities of the two incoherent optical sources are known and equal. This assumption looks trivial at first glance, but turns out to play a critical role in the superresolution: if the photon numbers are unknown and need to be estimated simultaneously with the separation of the two sources, the superresolution can work only when the intensities are equal and the estimation precision of the separation will vanish otherwise [85]. As the intensities could vary or fluctuate due to noises and technical imperfections in reality, this imposes challenges to the application of superresolution in practice. In recent years, the role of partial coherence in the superresolution has also been investigated intensively [86, 87, 88, 89, 90, 75, 91, 92], and entanglement has been found useful to enhance the precision of superresolution [93, 94].
In this work, we analyze the effect of unknown photon numbers on the precision of superresolution from the viewpoint of quantum multiparameter estimation in detail, and obtain the constraint condition on the difference of the photon numbers between the two incoherent point sources to avoid the vanishing of estimation precision. The results show the precision may deviate from that with identical photon numbers, depending on the magnitude of the difference of the photon numbers. We further consider two incoherent point sources with different point-spread functions, and find that the competition between the difference of the point-spread functions and the difference of the photon numbers determines the precision of superresolution for the separation. We analyze the competition between these parameters in detail, and show that the superresolution can have various precision limits depending on the magnitudes of these parameters, beyond the precision limit of two optical sources with identical photon numbers and point-spread functions.
II Preliminaries
Suppose two incoherent optical pointlike sources are located at unknown positions $X_{1}$ and $X_{2}$ on the one-dimensional object plane. Neglecting the probability that more than one photon simultaneously arrives at the image plane in the Poisson limit, the imaging process can be regarded as the events of detecting a series of single photons over multiple coherence time intervals of the sources [95]. Assume the intensities of the two optical sources can be different. Denote the average numbers of photons emitted by the two sources within a given observation time as $N_{1}$ and $N_{2}$ respectively. The average single-photon state of the optical field arriving at the image plane can be described by a density operator $\hat{\rho}$ [96],
$\hat{\rho}=\frac{1-\epsilon}{2}{\rm e}^{-iX_{1}\hat{P}}|\psi\rangle\langle\psi|{\rm e}^{iX_{1}\hat{P}}+\frac{1+\epsilon}{2}{\rm e}^{-iX_{2}\hat{P}}|\psi\rangle\langle\psi|{\rm e}^{iX_{2}\hat{P}}.$ (1)
The state $|\psi\rangle$ is defined as
[…]
where $\psi(x)$ is the point-spread functions of the imaging system assumed to be real and symmetric about the origin point [97] throughout this paper, $\epsilon$ is the ratio between the difference and the total of the photon numbers of the two optical sources,
$\epsilon=\frac{N_{2}-N_{1}}{N_{2}+N_{1}},$ (3)
and ${\rm e}^{-iX_{i}\hat{P}}$, $i=1,2$, are the displacement operators with $\hat{P}$ defined as
$\hat{P}=-i\partial_{x}.$ (4)
The probability distribution of a single photon arriving on the image plane is [20]
$\Lambda(x)=\frac{1-\epsilon}{2}\left|\langle x|{\rm e}^{-iX_{1}\hat{P}}|\psi\rangle\right|^{2}+\frac{1+\epsilon}{2}\left|\langle x|{\rm e}^{-iX_{2}\hat{P}}|\psi\rangle\right|^{2},$ (5)
and the photon count at a width $dx$ around the position $x$ of the detector is approximately $(N_{1}+N_{2})\Lambda(x)dx$.
Now, the question is: how precise can one measure and estimate the separation,
$d=X_{2}-X_{1},$ (6)
between the two incoherent point sources?
The well-known Rayleigh’s criterion in optical imaging and the quantum superresolution scheme proposed in the seminal paper [20] concern the distinguishability of two pointlike optical sources with identical point-spread functions up to a small spatial translation, for which the quantum model is given by Eq. (1). In this paper, we will revisit the quantum superresolution scheme for two point sources with identical point-spread functions to investigate the validity condition of quantum superresolution, and extend it to the point sources with different point-spread functions in Sec. IV.
II.1 Direct imaging and Rayleigh’s curse
The most straightforward method to measure the separation is directly detecting and counting photons arrived on the image plane from the two point sources [98].
In order to quantify the precision limit of estimating the separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which is reviewed in Appendix A. For the task of resolving two point sources, the positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ of the two sources, are unknown, so $\bar{X}$ and $d$ are the parameters to measure and estimate.
The Cramér-Rao bound demonstrates that over all possible estimation strategies, the covariance matrix $C$ for the estimates of the parameters $\bar{X}$ and $d$ is lower bounded by
$C[\bar{X},d]\geq N_{{\rm tot}}^{-1}\mathcal{J}^{-1}[\bar{X},d],$ (7)
where $N_{{\rm tot}}$ is total number of photons from the two optical sources within a given detection time,
$N_{{\rm tot}}=N_{1}+N_{2},$ (8)
the $\geq$” sign represents semi-definite positivity of matrix, and $\mathcal{J}[\bar{X},d]$ denotes the Fisher information matrix,
[…]
Note that the Cramér-Rao bound can always be asymptotically saturated by the maximum likelihood estimation strategy when the total number of photons $N_{{\rm tot}}$ is sufficiently large [15].
For the direct imaging with an ideal continuum photodetector on the image plane, if the intensities and the point-spread functions of the two optical sources are the same, i.e., $N_{1}=N_{2}=N$, the probability distribution of a single photon arrived on the image plane can be written as
$\Lambda(x)=\frac{1}{2}I(x-\bar{X}+d/2)+\frac{1}{2}I(x-\bar{X}-d/2),$ (10)
where $I(x)$ is defined as
$I(x)=|\psi(x)|^{2}.$ (11)
When the separation $d$ is small, it can be shown [72] by the Cramér-Rao bound (7) that the estimation precision of the separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{8}d^{2}\alpha_{2}N,$ (12)
where
[…]
The detail of derivation is provided in Appendix A. Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the separation by direct imaging when the two point sources are sufficiently close to each other. This is the Rayleigh’s curse for direct imaging of two identical optical point sources [20].
The above result can be generalized to the case that the two point-spread functions are the same but the intensities are unequal. In this case, the probability distribution of a single photon on the image plane can be written as
$\Lambda(x)=\frac{1-\epsilon}{2}I(x-\bar{X}+d/2)+\frac{1+\epsilon}{2}I(x-\bar{X}-d/2).$ (14)
The estimation precision of the separation by direct imaging can be found [72] to be
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{d^{2}\alpha_{2}N_{\text{tot}}}{16}\left(1-\epsilon^{2}\right)^{2}.$ (15)
(See Appendix A for detail.) This indicates that one cannot get any information about the separation $d$ by direct imaging in the limit $d\rightarrow 0$ in the generalized case either.
II.2 Quantum-limited superresolution
To surpass the Rayleigh’s criterion, one needs to further resort to the quantum estimation theory, which optimizes the Cramér-Rao bound over all possible quantum measurement schemes, including generalized measurements that are described by Kraus operators [99].
For the current quantum superresolution problem, the positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate.
The quantum estimation theory for the superresolution problem is briefly reviewed in Appendix B, which tells that the estimation precision of $\bar{X}$ and $d$ is lower bounded by the following quantum Cramér-Rao’s inequality,
$C[\bar{X},\,d]\geq N_{{\rm tot}}^{-1}\mathcal{Q}^{-1}[\bar{X},\,d],$ (16)
where $C[\bar{X},\,d]$ is the covariance matrix for the estimates of $\bar{X},\,d$, $\mathcal{Q}[\bar{X},\,d]$ is the quantum Fisher information matrix about $\bar{X},\,d$ with entries
$\mathcal{Q}_{ij}(\hat{\rho})=\frac{1}{2}{\rm Tr}[\hat{\rho}\{\hat{\mathcal{L}}_{i},\hat{\mathcal{L}}_{j}\}],i,j=1,2,$ (17)
and ^$\hat{\mathcal{L}}_{i}$, ^$\hat{\mathcal{L}}_{j}$ are the symmetric logarithmic derivatives (SLD) of the density operator $\hat{\rho}$,
$\frac{1}{2}(\hat{\mathcal{L}}_{i}\hat{\rho}+\hat{\rho}\hat{\mathcal{L}}_{i})=\partial_{g_{i}}\hat{\rho},\;g_{i}=\bar{X},d.$ (18)
Note that the quantum Cramér-Rao bound for multi-parameter estimation is not always achievable due to the potential noncommutativity between the optimal measurements for different parameters. But if one is interested in the overall estimation precision of different parameters, a weight matrix $W$ can be imposed on the unknown parameters and the matrix quantum Cramér-Rao’s bound can be turned to a scalar form,
${\rm Tr}(WC)\geq N_{{\rm tot}}^{-1}{\rm Tr}(W\mathcal{Q}^{-1}).$ (19)
The saturation condition for the above scalar quantum Cramér-Rao bound is
${\rm Tr}(\hat{\rho}[\hat{\mathcal{L}}_{i},\hat{\mathcal{L}}_{j}])={\rm ImTr}(\hat{\rho}\hat{\mathcal{L}}_{i}\hat{\mathcal{L}}_{j})=0,$ (20)
by recognizing that the quantum Cramér-Rao bound is equivalent to the Holevo bound when this condition is satisfied and the latter is always achievable asymptotically with a sufficiently large number of states [100, 101].
For the current problem of quantum superresolution, the interested parameter is the separation $d$, so one can choose
$W=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},$ (21)
assuming $g_{1}=\bar{X}$ and $g_{2}=d$, and ${\rm Tr}(WC)$ gives the estimation variance of the separation. And the above saturation condition can indeed be satisfied by the current problem, due to the reality of $\psi(x)$ which makes the symmetric logarithmic derivatives ^$\hat{\mathcal{L}}_{i}$ and ^$\hat{\mathcal{L}}_{j}$ also real according to Eq. (18) and thus the compatibility condition ${\rm ImTr}(\hat{\rho}\hat{\mathcal{L}}_{i}\hat{\mathcal{L}}_{j})=0$ can be satisfied. So the quantum Cramér-Rao bound provides an achievable precision limit for the estimation of the separation allowed by quantum mechanics.
A striking result found by Tsang et al. [20] is that the best estimation precision of the separation $d$ over all quantum measurements for two incoherent point sources with the same photon numbers, determined by the quantum Cramér-Rao’s bound, keeps nonzero even when the separation goes to zero, implying the Rayleigh’s criterion is violated when the quantum measurement is optimized. The optimal precision of estimating the separation $d$ in the limit $d\rightarrow 0$ is found to be
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=2N\langle\hat{P}{}^{2}\rangle,$ (22)
[…]
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=N_{{\rm tot}}\left(1-\epsilon^{2}\right)\langle\hat{P}{}^{2}\rangle.$ (23)
For two Gaussian point-spread functions with the same width $\sigma$, the above precision limit can be simplified to
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)}{4\sigma^{2}}.$ (24)
The derivation of the precision limits for a general $d$ and for the above $d\rightarrow 0$ case is presented in Appendix C.1.
An interesting finding in recent years is the vanishing of estimation precision when the photon numbers $N_{1}$, $N_{2}$, of two point sources are unknown and need to be estimated as well [85]. In this case, one needs to estimate three parameters, $\bar{X},\,d,\,\epsilon$, simultaneously. If the photon numbers of the two point sources are different, i.e., $\epsilon\neq 0$, it was found that up to the second order of $d$, the precision of estimating $d$ was
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)}{4\epsilon^{2}}{\rm Var}(\hat{P}^{2})d^{2},$ (25)
where ${\rm Var}(\hat{P}^{2})=\langle\hat{P}^{4}\rangle-\langle\hat{P}^{2}\rangle^{2}$.
It can be immediately seen that in this case, $\mathcal{H}_{d}$ is of order $O(d^{2})$ and will vanish as $d$ goes to zero, implying the superresolution no longer works and the Rayleigh’s curse applies again. This is sometimes termed as the resurgence of Rayleigh’s curse for two incoherent point sources with unknown and unbalanced intensities [85, 102].
III Superresolution for two identical point sources
In this work, we investigate the problem of precision vanishing in resolving two incoherent point sources with unknown and unequal intensities and study the condition for the quantum superresolution to address this problem in detail.
As a finite difference between the photon numbers of the two sources can cause the resurgence of Rayleigh’s curse while the equality between the two photon numbers leads to superresolution, we are interested in the condition on the photon numbers to realize the superresolution: is it possible to realize the superresolution with a sufficiently small but nonzero difference between the photon numbers of the two sources and how close the two photon numbers need to be? In addition to the photon numbers, we will also investigate the effect of slight difference between the point-spread functions of the two optical sources on suppressing the Rayleigh’s criterion, and show its competition with the difference of the photon numbers in determining the precision of resolving the two point sources.
We consider two identical incoherent point sources located at $X_{1}$, $X_{2}$ with unknown photon numbers $N_{1}$, $N_{2}$ in this section. The unknown parameters to estimate in this case is $\bar{X}$, $d$, $\epsilon$, and $\epsilon$ is the ratio between the difference and the total of the photon numbers defined in Eq. (3). The density matrix $\hat{\rho}$ of a single photon is given by Eq. (1). By working out the eigenvalues and eigenstates of $\hat{\rho}$ and applying the formula of the quantum Fisher information matrix, the quantum Fisher information matrix with respect to these three unknown parameters can be obtained [85] as
$\mathcal{Q}=\left[\begin{array}[]{ccc}4\left[\kappa-\gamma^{2}\left(1-\epsilon^{2}\right)\right]&2\kappa\epsilon&2\gamma\delta\\ 2\kappa\epsilon&\kappa&0\\ 2\gamma\delta&0&\frac{1-\delta^{2}}{1-\epsilon^{2}}\end{array}\right],$ (26)
where
$\kappa=\langle\hat{P}{}^{2}\rangle,\;\delta=\langle\cos(d\hat{P})\rangle,\;\gamma=\langle\sin(d\hat{P})\hat{P}\rangle.$ (27)
According to the quantum Cramér-Rao bound, the estimation precision of the separation $d$ can be obtained as
$\mathcal{H}_{d}=N_{{\rm tot}}/(\mathcal{Q}^{-1})_{22}=\frac{\left(1-\epsilon^{2}\right)\left[\gamma^{2}-\left(1-\delta^{2}\right)\kappa\right]}{\gamma^{2}\left(1-\epsilon^{2}\right)-(1-\delta^{2})\kappa}\mathcal{H}_{d}^{({\rm opt})},$ (28)
where $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision when the photon numbers of the two sources are equal, i.e., $\epsilon=0$, $N_{1}=N_{2}=N$,
$\mathcal{H}_{d}^{({\rm opt})}=2\kappa N,$ (29)
The detail of derivation is provided in Appendix C.2.
To investigate the mechanism underlying the vanishing of estimation precision when the photon numbers $N_{1}$ and $N_{2}$ of the two sources are different, we expand the numerator and denominator of the precision $\mathcal{H}_{d}$ (28) to the fourth order of $d$ respectively, and the result turns out to be
$\mathcal{H}_{d}=\frac{Cd^{2}}{D_{1}+D_{2}d^{2}},$ (30)
where
$C=$ $3N_{{\rm tot}}(1-\epsilon^{2})\langle\hat{P}{}^{2}\rangle{\rm Var}(\hat{P}^{2}),$ (31)
$D_{1}=$ $12\epsilon^{2}\langle\hat{P}{}^{2}\rangle,$
$D_{2}=$ $3{\rm Var}(\hat{P}^{2})-4\epsilon^{2}\langle\hat{P}^{4}\rangle,$
and a common $d^{2}$ has been reduced from both the numerator and denominator.
Obviously, when the intensities of the two optical sources are unequal, i.e., $N_{1}\neq N_{2}$, the coefficient $D_{1}$ is finite, so $\mathcal{H}_{d}=O(d^{2})$, thus $\mathcal{H}_{d}$ vanishes when $d\rightarrow 0$, and the Rayleigh’s curse takes effect again. In contrast, when the intensities of the two optical sources are equal, i.e., $\epsilon\rightarrow 0$ and $N_{{\rm tot}}=N_{1}+N_{2}=2N$, $D_{1}=0$, so the estimation precision of $d$ (30) becomes
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{C}{D_{2}}=2N\langle\hat{P}{}^{2}\rangle,$ (32)
and the superresolution scheme can work in this case, which is the result by Ref. [20].
Nevertheless, a further observation of $\mathcal{H}_{d}$ in Eq. (30) suggests that even when $N_{1}\neq N_{2}$ there is still opportunity to have nonvanishing estimation precision for the separation $d$ if the coefficient $D_{1}$ is sufficiently small. It can be seen that if $D_{1}$ has the order $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$, i.e., the order of $\epsilon$ is $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, the precision $\mathcal{H}_{d}$ keeps finite when $d$ is small, implying the superresolution scheme can still work in this case. Furthermore, if the order of the coefficient $D_{1}$ is higher than $O(d^{2}{\rm Var}(\hat{P}^{2}))$, or equivalently the order of $\epsilon$ is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, $D_{1}$ can be dropped when $d\rightarrow 0$, and $\mathcal{H}_{d}$ is equal to $C/D_{2}=N_{{\rm tot}}\langle\hat{P}{}^{2}\rangle$ which restores the original result for the precision of superresolution in [20].
The above observation immediately leads to the condition for the superresolution of two identical incoherent optical sources: the order of the ratio $\epsilon$ between the difference and the total of the photon numbers of the two optical sources must be higher than or equal to $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$. The original result in [20] was obtained for the case $N_{1}=N_{2}$, but is effective for the range of $\epsilon$ that is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, while the superresolution scheme can actually work in a broader range including the case that the order of $D_{1}$ is equal to $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$. But the latter has a lower precision than the former, as the ratio between the two precisions is
$\frac{D_{2}d^{2}}{D_{1}+D_{2}d^{2}}<1.$ (33)
It must be noted that this condition for the validity of superresolution with unknown photon numbers does not mean that the photon number changes with $d$, but rather impose a limitation on the magnitude of photon number difference between the two optical sources to realize the superresolution.
IV Superresolution for two arbitrary point sources
While the point-spread functions can be identical for the point sources at different positions of the object plane if the imaging system is spatially invariant with specific conditions such as the paraxial approximation [20, 27, 103], the point-spread functions can vary with the positions of the point sources in general [104]. For example, the longitudinal shifts from the focused object plane will make the locations of two point sources deviate in the depth from each other, and thus lead to different point-spread functions due to the diffraction and geometrical blurring effects [105, 106, 107]. Such a difference in the point-spread functions may reduce the overlap between the two optical sources, so an intuitive idea is that it may increase the distinguishability of the two sources which could be helpful to the resolution of the two sources. In the following, we generalize the result of the previous section and study the condition of quantum superresolution for two optical point sources with arbitrary point-spread functions.
[…]
$\hat{\rho}=$ $\frac{1-\epsilon}{2}{\rm e}^{-iX_{1}\hat{P}}|\psi_{1}\rangle\langle\psi_{1}|{\rm e}^{iX_{1}\hat{P}}$ (34)
$+\frac{1+\epsilon}{2}{\rm e}^{-iX_{2}\hat{P}}|\psi_{2}\rangle\langle\psi_{2}|{\rm e}^{iX_{2}\hat{P}},$
where $N_{{\rm tot}}$ is the total number of the photons, $N_{{\rm tot}}=N_{1}+N_{2}$, and $\epsilon$ is the ratio between the difference and the total of the photon numbers of the two point sources defined in Eq. (3).
The estimation precision for the separation can still be obtained by the scalar quantum Cramér-Rao bound (19). The quantum Fisher information matrix with respect to the unknown parameters $\bar{X},d,\epsilon$ turns out to be
$\mathcal{Q}=$ (35)
$\left[\begin{array}[]{ccc}2\kappa_{\text{tot}}(1+\chi\epsilon)-4\gamma^{2}\left(1-\epsilon^{2}\right)&\kappa_{\text{tot}}(\chi+\epsilon)&2\gamma\delta\\ \kappa_{\text{tot}}(\chi+\epsilon)&\frac{1}{2}\kappa_{\text{tot}}(1+\chi\epsilon)&0\\ 2\gamma\delta&0&\frac{1-\delta^{2}}{1-\epsilon^{2}}\end{array}\right],$
leading to the exact result for the estimation precision of $d$ as
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)\left[\left(1-\delta^{2}\right)\left(1-\chi^{2}\right)\kappa_{{\rm tot}}-2\gamma^{2}(1+\chi\epsilon)\right]}{\left(1-\delta^{2}\right)(1+\chi\epsilon)\kappa_{{\rm tot}}-2\gamma^{2}\left(1-\epsilon^{2}\right)}\mathcal{H}_{d}^{({\rm opt})},$ (36)
where
$\kappa_{i}=$ $\langle\hat{P}{}^{2}\rangle_{i},\,i=1,2,$ (37)
$\delta=$ $\langle\cos(d\hat{P})\rangle_{12},\;\gamma=\langle\sin(d\hat{P})\hat{P}\rangle_{12},$
$\kappa_{{\rm tot}}=$ $\langle\hat{P}{}^{2}\rangle_{1}+\langle\hat{P}{}^{2}\rangle_{2},\;\chi=\frac{\langle\hat{P}{}^{2}\rangle_{2}-\langle\hat{P}{}^{2}\rangle_{1}}{\langle\hat{P}{}^{2}\rangle_{1}+\langle\hat{P}{}^{2}\rangle_{2}},$
and $\langle\cdot\rangle_{i}=\langle\psi_{i}|\cdot|\psi_{i}\rangle$, $\langle\cdot\rangle_{12}=\langle\psi_{1}|\cdot|\psi_{2}\rangle$. $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision when the photon numbers of the two sources are equal, i.e., $\epsilon=0$, in the limit $d\rightarrow 0$,
$\mathcal{H}_{d}^{({\rm opt})}=\frac{1}{2}N_{{\rm tot}}\kappa_{{\rm tot}},$ (38)
which can be simplified to Eq. (29) when the point-spread functions of the two sources are the same. The derivation is provided in Appendix D.
Considering the separation $d$ is small, we expand the numerator and the denominator of $\mathcal{H}_{d}$ to the fourth order of $d$ respectively,
$\mathcal{H}_{d}=\frac{A_{0}+A_{2}d^{2}+A_{4}d^{4}}{B_{0}+B_{2}d^{2}+B_{4}d^{4}}.$ (39)
The coefficients $A_{i}$, $B_{i}$, $i=0,2,4$, are quite complex, so they are provided in Eq. (126) of Appendix D.
It is obtained in Appendix D that
$A_{0},B_{0}\propto 1-|\langle\psi_{1}|\psi_{2}\rangle|^{2}.$ (40)
So, if the difference between the two point-spread functions is not small, i.e., the magnitudes of $A_{0}$ and $B_{0}$ are of order $O(1)$, only the constant terms $A_{0}$ and $B_{0}$ remain in $\mathcal{H}_{d}$ when $d\rightarrow 0$. The estimation precision of the separation $d$ in the limit $d\rightarrow 0$ can be simplified to
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{N_{{\rm tot}}\kappa_{{\rm tot}}\left(1-\chi^{2}\right)\left(1-\epsilon^{2}\right)}{2(1+\chi\epsilon)},$ (41)
implying that the superresolution can work even if the photon numbers of the two point sources are different in this case.
A more subtle case is when the point-spread functions of the two optical sources are different but close to each other. In this case, the constant terms $A_{0}$ and $B_{0}$ in $\mathcal{H}_{d}$ (39) will become small though still nonzero. If the magnitudes of these leading terms are comparable to $O(d^{2})$ or even $O(d^{4})$ (although they are independent of $d$ in physics), they will compete with the second-order and fourth-order terms in Eq. (39) to determine $\mathcal{H}_{d}$.
Similarly, it is shown in Appendix D that the coefficients $A_{2}$ and $B_{2}$ of the quadratic terms in $\mathcal{H}_{d}$ may also become small but nonzero when the two point-distribution functions differ slightly from each other, so if they are of order $O(d^{2})$, they will also compete with the fourth-order terms in determining $\mathcal{H}_{d}$. Moreover, the difference between the photon numbers of the two optical sources are also involved in $\mathcal{H}_{d}$ which may compete with the higher-order terms in $\mathcal{H}_{d}$ if the difference in the photon numbers of the two point sources is small. Combining all these together, the estimation precision of the separation $d$ and thus the validity of superresolution are highly dependent on the competition between the difference in the two point-spread functions, the difference in the photon numbers of the two point sources and the magnitude of the separation.
Table 1: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ between the estimation precision $\mathcal{H}_{d}$ of the separation $d$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (38) in different regimes of the parameters. The difference between the point-spread functions and that between the photon numbers are characterized by Eq. (44) and we denote $1-\langle\psi_{1}|\psi_{2}\rangle=b\text{$\widetilde{d}^{h}$}$, $\frac{1}{2}-\xi=a\widetilde{d}^{f}$, $\chi=c\widetilde{d}^{e}$, $\epsilon=y\widetilde{d}^{s}$, where $y,a,b,c$ are real coefficients, and $\widetilde{d}$ is the dimensionless separation (43). The exponents $s,f,h,e$ are nonnegative integers, and $h,\,f,\,e$ satisfy the constraint relation (45). The blue region represents the cases that attain the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region represents the cases that the estimation precision vanishes, and the yellow region represents the cases that the superresolution can still work but with a lower precision. The estimation precision of the separation $d$ given by the original superresolution scheme in [20] corresponds to $h\geq 3$, $s\geq 2$, and the resurgence of Rayleigh’s curse [85] corresponds to $h\geq 3,\,s=0$. The total photon number $N_{{\rm tot}}$ is sufficiently large, i.e., $N_{{\rm tot}}\gg 1$. The results are evaluated in the limit $d\rightarrow 0$. As the two point-spread functions approach to be the same in this limit, the subscripts of $\left\langle P^{4}\right\rangle{}_{12}$ and $\left\langle P^{2}\right\rangle{}_{12}$ are dropped in the table.
$\ s\geq 2\\ \\ \end{gathered}$ $\ s=1\\ \\ \end{gathered}$ $\ s=0\\ \\ \end{gathered}$
$\ h\geq 5\\ \\ \end{gathered}$ $\ \frac{{\rm Var}(\hat{P}^{2})}{4y^{2}\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$
$\ h=4\\ \\ \end{gathered}$ $\ \frac{8b\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}{4(2b+y^{2})\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$ $\begin{gathered}0\\ \\ \\ \end{gathered}$
$\ h=3\\ \\ \end{gathered}$
$\ h=2\\ \\ \end{gathered}$ $\ 1\\ \\ \end{gathered}$ $\begin{gathered}\frac{2b\left(1-y^{2}\right)}{2b+y^{2}}\end{gathered}$
$\ h=1\\ \\ \end{gathered}$ $\begin{gathered}1-y^{2}\end{gathered}$
$\ h=0\\ \\ \end{gathered}$ $\begin{gathered}1-c^{2}\end{gathered}$ $\begin{gathered}\frac{\left(1-c^{2}\right)\left(1-y^{2}\right)}{1+cy}\end{gathered}$
As the point-spread functions are continuous, it generally requires an infinite number of parameters to characterize the difference between the two point-spread functions in principle, which seems to impose significant difficulty in analyzing the effect of the difference of the point-spread functions on the precision of quantum superresolution. But in fact, one can figure out just a few quantities from $\mathcal{H}_{d}$ (39) that are dependent on these differences and will determine $\mathcal{H}_{d}$. In Appendix E.1, it is shown that only the following three quantities, $1-\langle\psi_{1}|\psi_{2}\rangle$, $\chi$ as defined in Eq. (37), and $\frac{1}{2}-\xi$ with $\xi$ defined as
$\xi=\frac{\langle\hat{P}{}^{2}\rangle_{12}}{\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}},$ (42)
are relevant to the difference between the two point-spread functions and will determine the estimation precision $\mathcal{H}_{d}$. And they tend to be zero as the point-spread functions of the two sources get close to each other, so these quantities will compete with each other and with the difference of photon numbers and play a key role in determining the validity of superresolution.
But note that the photon numbers and the closeness between the point-spread functions are dimensionless while the separation $d$ is not, so for the sake of convenience to make these quantities comparable and characterize the magnitudes of these quantities, we rescale the separation $d$ to make it dimensionless. We define the following dimensionless separation $\widetilde{d}$,
$\widetilde{d}=\frac{d}{V},$ (43)
where $V=1/\sqrt{\langle\hat{P}{}^{2}\rangle}$. The subscript is dropped from $\langle\hat{P}{}^{2}\rangle$ in the definition of $V$, as the following results of the estimation precision $\mathcal{H}_{d}$ are evaluated in the limit $d\rightarrow 0$ which leads to $\psi_{1}(x)=\psi_{2}(x)$ (as the difference between the two point-spread functions is characterized in terms of the power of $\widetilde{d}$ later) and thus $\langle\hat{P}{}^{2}\rangle_{1}=\langle\hat{P}{}^{2}\rangle_{2}=\langle\hat{P}{}^{2}\rangle_{12}$. In order to represent the magnitudes of these quantities which are all considered to be small, we denote them in the orders of $\widetilde{d}$,
$1-\langle\psi_{1}|\psi_{2}\rangle=O(\widetilde{d}^{h}),$ $\frac{1}{2}-\xi=O(\widetilde{d}^{f}),$ (44)
$\epsilon=O(\widetilde{d}^{s}),$ $\chi=O(\widetilde{d}^{e}),$
where $s,f,h,e$ are nonnegative integers. Note that the exponents $h,f,e$ cannot be arbitrary but are constrained by the relation
$h/2\leq e\leq h\leq f\leq 2h.$ (45)
A detailed proof of this relation can be found in Appendix E.2. This constraint will be taken into account in the following computation of $\mathcal{H}_{d}$. When the exponents $h,f,e,s$ vary, the estimation precision $\mathcal{H}_{d}$ of the separation $d$ will be different. The estimation precision $\mathcal{H}_{d}$ can be computed by Eq. (39) for different ranges of these exponents.
The results of the estimation precision $\mathcal{H}_{d}$ are summarized in Table 1 in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (38), which covers all the possible ranges of the difference between the two point-spread functions and the difference between the two photon numbers and rules out the cases that the exponents do not satisfy the constraint relation (45). It can be seen that the superresolution protocol can actually work for a wide range of point-spread functions with various magnitudes of the photon numbers, and the estimation precision $\mathcal{H}_{d}$ vanishes only when the order of $1-\langle\psi_{1}|\psi_{2}\rangle$ is higher than or equal to $O(\widetilde{d}^{3})$ and $\epsilon=O(1)$, which is the cases considered in Ref. [85]. The case $h\geq 3$ with $s\geq 2$ is the regime that was considered by the original superresolution scheme [20]. The case $h\geq 5$ with $s=1$ is the regime where the effective condition of quantum superresolution for two identical point-spread functions was obtained in Sec. III. For the other cases, the superresolution scheme can still work, but the estimation precision of the separation $d$ may deviate from that of the original superresolution scheme [20].
It needs to be stressed again that the orders of the difference between the point-spread functions and the difference between the photon numbers of the two point sources in terms of the (dimensionless) separation do not mean that they change with $d$ but rather impose a limitation on the magnitudes of those differences to realize the superresolution.
V Example: Point sources of Gaussian distributions
To illustrate the above results, we consider the Gaussian point-spread functions as a typical example. Assume $\psi_{1}(x)$ and $\psi_{2}(x)$ are Gaussian point-spread functions,
$\psi_{i}(x)=\frac{1}{\sqrt[4]{2\pi\sigma_{i}^{2}}}\exp\left[-\frac{x^{2}}{4\sigma_{i}^{2}}\right],\;i=1,2,$ (46)
where $\sigma_{i}=\frac{1}{2\sqrt{\kappa_{i}}}=\lambda_{i}/2\pi{\rm NA},i=1,2$, $\lambda$ is the free-space wavelength, ${\rm NA}$ is the effective numerical aperture [108], and the spatial-frequency variance $\kappa_{i}$ for the Gaussian point-spread function is $\frac{1}{4\sigma_{i}^{2}},\,i=1,2$. The two general Gaussian point-spread functions are illustrated in Fig. 1.
V.1 Gaussian point-spread functions with the same width
For two Gaussian point-spread functions with the same width $\sigma$, the estimation precision of the separation is given by Eq. (30), where the coefficients can be simplified by the Gaussian point-spread function and the estimation precision of $d$ becomes
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)d^{2}/\sigma^{2}}{32\epsilon^{2}+4d^{2}/\sigma^{2}}\mathcal{H}_{d}^{({\rm opt})},$ (47)
where $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision of $d$ when the photon numbers of the two point sources are equal and $d\rightarrow 0$,
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N_{{\rm tot}}}{\sigma^{2}}.$ (48)
It can be seen that the ratio between $\mathcal{H}_{d}$ and $\mathcal{H}_{d}^{({\rm opt})}$ is determined by two parameters only, $\epsilon$, i.e., the ratio between the difference and the total of the two photon numbers, and $d/\sigma$, i.e., the ratio between the separation and the width of the two point sources.
Figure 1: Illustration of two Gaussian point-spread functions $\psi_{1}(x)$ and $\psi_{2}(x)$ with positions $X_{1}$, $X_{2}$ and widths $\sigma_{1}$, $\sigma_{2}$ respectively. The purpose of the superresolution scheme is to estimate the separation $d=X_{2}-X_{1}$ when $d$ is small.
Hence, if the difference between the two photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$, the estimation precision $\mathcal{H}_{d}$ vanishes as $d$ goes to 00, implying the Rayleigh’s curse applies again. But if the order of the ratio $\epsilon$ is higher than or equal to $O(d/\sigma)$, the magnitude of the denominator of $\mathcal{H}_{d}$ becomes comparable to the numerator, so the estimation precision $\mathcal{H}_{d}$ will be independent of $d$, and one can still acquire information about the separation $d$ in this case. In particular, if the order of $\epsilon$ is higher than $O(d/\sigma)$, $\mathcal{H}_{d}$ will be restored to the original estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with equal photon numbers obtained in [20]. If the order of $\epsilon$ is equal to $O(d/\sigma)$, the precision $\mathcal{H}_{d}$ does not vanish, but will be decreased by a factor (33),
$\frac{1-\epsilon^{2}}{8\epsilon^{2}\left(\frac{\sigma}{d}\right)^{2}+1}.$ (49)
This explicitly shows the influence of the photon number difference in the estimation precision of superresolution.
To illustrate the above result numerically, the ratio of the precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (48) for two Gaussian point-spread functions with the same width is plotted with respect to $\epsilon$ for different $d/\sigma$ in Fig. 2. It can be seen that for a fixed separation, the precision of the separation reaches the maximum when $\epsilon=0$, i.e., the photon numbers of the two point sources are equal, which is the regime of the original superresolution in [20], and decreases to zero when the photon number difference increases, which is the reoccurrence of Rayleigh’s curse found in [85]. It also shows that when $\epsilon$ is nonzero but sufficiently small, the precision $\mathcal{H}_{d}$ will still be significant and the superresolution can work, showing the effective range of quantum superresolution which is broader than $\epsilon=0$. Moreover, the figure also suggests that for the same $\epsilon$, a smaller separation compared to the width of the Gaussian point-spread function leads to a lower estimation precision for $d$. This can be understood as for a smaller $d$, the magnitude of $\epsilon$ becomes larger compared to $d/\sigma$, so it is more difficult to satisfy the effective condition of the superresolution and the estimation precision of $d$ decreases as a result.
V.2 Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$
For two Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$, the constant terms in Eq. (30) can be obtained as
$A_{0}=$ $8N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left(\sigma_{1}-\sigma_{2}\right){}^{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2},$ (50)
$B_{0}=$ $16\left(\sigma_{1}-\sigma_{2}\right){}^{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right].$
Figure 2: Plot of the ratio of the estimation precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ for two Gaussian point-spread functions given different ratios $d/\sigma$. $d$ is the separation of the two point sources, and $\sigma$ is the width of the two Gaussian point-spread functions. It can be observed from the figure that when the photon numbers of the two sources are the same, i.e., $\epsilon=0$, the estimation precision can always reach the optimum, and when $\epsilon>0$, the estimation precision decreases. The ratio $d/\sigma$ influences how fast the precision $\mathcal{H}_{d}$ decays to zero, i.e., the speed that $\mathcal{H}_{d}$ transits from the regime of quantum superresolution to the regime of Rayleigh’s curse, when the photon number difference increases. It verifies Eq. (62) that $d/\sigma=0$ is the minimal point of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$, so the large the ratio $d/\sigma$ is, the higher the precision $\mathcal{H}_{d}$ is and hence the slower $\mathcal{H}_{d}$ decays to zero, and vice versa.
If we take the limit $d\rightarrow 0$ immediately in $\mathcal{H}_{d}$ (39), the estimation precision $\mathcal{H}_{d}$ turns out to be
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{1-\epsilon^{2}}{1-2\eta\epsilon+\eta^{2}}\mathcal{H}_{d}^{({\rm opt})},$ (51)
where
$\eta=\frac{\sigma_{2}-\sigma_{1}}{\sigma_{1}+\sigma_{2}},\;\sigma_{{\rm tot}}=\sigma_{1}+\sigma_{2},$ (52)
and $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision when the widths of the two Gaussian point-spread functions are the same, $\sigma_{1}=\sigma_{2}=\sigma$ and the photon numbers of the two point-sources are equal, $N_{1}=N_{2}=N$,
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N_{{\rm tot}}}{\sigma_{{\rm tot}}^{2}}=\frac{N}{2\sigma^{2}}.$ (53)
Eq. (51) tells that the ratio $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is determined by the ration $\eta$ between the difference and the total of the widths of the two Gaussian point-spread functions and the ratio $\epsilon$ between the difference and the total of the two photon numbers only.
It is worth mentioning that there is a trade-off relation between the ratios $\epsilon$ and $\eta$ in the estimation precision $\mathcal{H}_{d}$ (51). By optimizing $\mathcal{H}_{d}$ (51) over all $\epsilon$ and $\eta$, it can found that the estimation of the separation $d$ reaches the highest precision when $\epsilon=\eta$, and the precision $\mathcal{H}_{d}$ is exactly the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (53) in this case. This implies that the optimal precision of estimating the separation $d$ can even be achieved when the photon numbers of the two point sources are different, if the widths of the two Gaussian point-spread functions are also different and the condition $\epsilon=\eta$ is satisfied.
It should be noted that such a result of $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ implicitly assumes that the order of the leading terms $A_{0}$ and $B_{0}$ in $\mathcal{H}_{d}$ (39) is $O(1)$, i.e., the width difference $\sigma_{2}-\sigma_{1}$ is sufficiently large compared to the separation $d$. When the two Gaussian optical sources are just slightly different in the widths, i.e., $\sigma_{2}-\sigma_{1}$ is small, higher-order terms of $d$ in $\mathcal{H}_{d}$ can also come into play in determining $\mathcal{H}_{d}$ when $d\rightarrow 0$. Similar to the preceding section, the precision $\mathcal{H}_{d}$ for the separation $d$ will be determined by the competition between the difference in the widths of Gaussian point-spread functions and the difference in the photon numbers in this case.
To characterize the magnitudes of the difference between the two widths and the difference between the two photon numbers, we denote the ratio between the difference and the total of the photon numbers and the ratio between the difference and the total of the two widths in the powers of $\tilde{d}$ as
$\epsilon=\frac{N_{2}-N_{1}}{N_{{\rm tot}}}=O(\widetilde{d}^{s}),\;\eta=\frac{\sigma_{2}-\sigma_{1}}{\sigma_{{\rm tot}}}=O(\widetilde{d}^{t}),$ (54)
so that we can compare the magnitudes of the photon number difference and the width difference and obtain the estimation precision $\mathcal{H}_{d}$ of the separation $d$ in different ranges of the these parameters. For the simplicity of computation, we redefine the dimensionless separation $\widetilde{d}$(43) as
$\widetilde{d}=\frac{d}{(\sigma_{1}+\sigma_{2})/2}.$ (55)
In the limit $d\rightarrow 0$, $\sigma_{1}=\sigma_{2}$ as the difference between the two point-spread functions is quantified in terms of the power of $\widetilde{d}$ (54), and $\widetilde{d}$ can now be simplified to
$\widetilde{d}=\frac{d}{\sigma}.$ (56)
Table 2: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ (53) for Gaussian point-spread functions with width difference $\eta=z\widetilde{d}^{t}$ and photon number difference $\epsilon=y\widetilde{d}^{s}$. The blue region denotes the cases that reach the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region denotes the cases that the estimation precision vanishes, and the yellow region denotes the cases that the superresolution work with a lower precision. The case $t\geq 2$, $s\geq 2$, where the magnitudes of the coefficients $z$ and $y$ are $O(1)$, includes the original superresolution scheme for two Gaussian point spread functions, and the case $t\geq 2$, $s=0$ is the resurgence of Rayleigh’s curse where the estimation precision of the separation vanishes. All the results are evaluated in the limit $d\rightarrow 0$, and the dimensionless separation $\widetilde{d}$ is given by Eq. (56).
$\ s\geq 2\\ \\ \end{gathered}$ $\ s=1\\ \\ \end{gathered}$ $\ s=0\\ \\ \end{gathered}$
[…]
$\ t=1\\ \\ \end{gathered}$ $\ \frac{8\left(1-y^{2}\right)z^{2}}{y^{2}+8z^{2}}\\ \\ \end{gathered}$
$\ t=0\\ \\ \end{gathered}$ $\ \frac{1}{1+z^{2}}\\ \\ \end{gathered}$ $\ \frac{1-y^{2}}{1-2yz+z^{2}}\\ \\ \end{gathered}$
Table 2 summarizes the results of $\mathcal{H}_{d}$ in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (53) for two arbitrary Gaussian point-spread functions in the limit $d\rightarrow 0$ for all feasible ranges of the exponents $t$ and $s$, which shows that different magnitudes of the difference in the photon numbers and the difference in the widths lead to different estimation precisions of the separation. Similar as the general results in the preceding section, the Rayleigh’s criterion applies only in the case that the widths of the two Gaussian point-distribution functions are sufficiently close, i.e., $t\geq 2$, and the difference of the photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$. For all the other cases, the superresolution scheme can work. The case $t\geq 2$, i.e., the ratio of the difference of the widths to the total of the widths is of order $O(\widetilde{d}^{2})$ or higher, with $\epsilon\leq O(\widetilde{d}^{2})$, includes the case considered by the original superresolution scheme [20] for two identical Gaussian optical sources with equal photon numbers. For the other regimes of the above two ratios, the quantum resolution protocol is still effective, but the estimation precision may vary. In particular, when $t\leq 1$, i.e., the order of the ratio between the difference and the total of the two Gaussian widths is lower than or equal to $O(\widetilde{d})$, the superresolution can always work regardless of the photon number difference, implying the difference between the two Gaussian point-spread functions is sufficient to make the superresolution work in this case, which is a manifestation of the effect of the difference between the point-spread functions in extending the effective regime of superresolution.
The superresolution precision $\mathcal{H}_{d}$ for two different Gaussian point-spread functions is plotted in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respective to $\epsilon$ in Fig. 3. It can be seen from Fig. 3a that given the separation $d=0$, for each $\eta$, the optimal estimation precision can always be achieved, but by different $\epsilon$. In fact, the estimation precision reaches the maximum when $\epsilon=\eta$, which can be obtained by optimizing Eq. (51) over $\epsilon$ and $\eta$ as discussed above. And the figure verifies that optimization result. Fig. 3b considers the case $\eta=0.005$ for different dimensionless separations $\widetilde{d}$ of the two Gaussian point-spread functions. It shows that the estimation can always reach the highest precision and the quantum superresolution works when the photon numbers of the two sources are equal, i.e., $\epsilon=0$. And as $\epsilon$ increases, the estimation precision $\mathcal{H}_{d}$ will decrease to zero and the Rayleigh’s curse takes effect again.
Moreover, one can see that the precision $\mathcal{H}_{d}$ does not change monotonically with the dimensionless separation $\widetilde{d}$ and the ratio $\eta$. In fact, there exist extreme points of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$ and $\eta$, when $\epsilon$ is fixed.
Figure 3: Plot of the ratios between the estimation precision $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ for two different Gaussian point-spread functions. (a) The estimation precision $\mathcal{H}_{d}$ for different ratios $\eta$ of the width difference to the total width of the two sources in the limit $d\rightarrow 0$. It can be observed that the precision $\mathcal{H}_{d}$ can still reach the maximum though the photon numbers of the two point sources are different if the point-spread functions of the two point sources are not identical, and the highest precision is achieved when $\epsilon=\eta$. The transition of $\mathcal{H}_{d}$ to the regime of Rayleigh’s curse starts when the difference of the photon numbers increases beyond the point $\epsilon=\eta$, and the larger $\eta$ is, the faster the estimation precision $\mathcal{H}_{d}$ decays with $\epsilon$. The exact effect of $\eta$ on the decay of the estimation precision is determined by Eq. (51). (b) The estimation precision $\mathcal{H}_{d}$ for different dimensionless separations $\widetilde{d}$ (55). The ratio $\eta$ (54) is $0.005$. The estimation precision can always reach the maximum when the photon numbers of the two sources are equal, i.e., $\epsilon=0$, and decays to zero as $\epsilon$ increases. The dimensionless separation $\widetilde{d}$ changes the decay speed of the estimation precision. $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ has two extreme points in this case: the maximal point is $\widetilde{d}=0$, and the minimal point is $\widetilde{d}\approx 0.2$, which can be determined by Eq. (57). So, when $0\leq\widetilde{d}\apprle 0.2$, $\mathcal{H}_{d}$ becomes smaller and decays faster with $\epsilon$ when $\widetilde{d}$ increases, and when $\widetilde{d}\apprge 0.2$, $\mathcal{H}_{d}$ behaves in the opposite way.
When $\eta\neq 0$, i.e., the widths of the two Gaussian point-spread functions are different, the extreme points of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$ can be worked out as
$\widetilde{d}_{1}=$ $0,$ (57)
$\widetilde{d}_{2}=$ $2\sqrt{\eta^{2}+1}\sqrt{w_{0}+1},$
where $w_{0}$ is the value of the principal branch $W_{0}$ of the Lambert W function at $\frac{\eta^{2}-1}{e\left(\eta^{2}+1\right)}$, i.e., $w_{0}=W_{0}\left[\frac{\eta^{2}-1}{e\left(\eta^{2}+1\right)}\right]$. When the ratio $\eta$ is small, we can expand the above second solution to $\widetilde{d}$ to the lowest order of $\eta$,
$\widetilde{d}_{2}\approx 2\sqrt{2\eta}.$ (58)
In order to determine the maximal point and the minimal point from the above two extremal solutions, we compute the second-order derivative of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$, and it turns out that
$\left.\partial_{\widetilde{d}}^{2}\mathcal{H}_{d}\right|_{\widetilde{d}=r_{1}}=-\frac{\left(1-\eta^{2}\right)\left(1-\epsilon^{2}\right)\left(-2\eta+\eta^{2}\epsilon+\epsilon\right)^{2}}{4\eta^{2}\left(\eta^{2}+1\right)^{2}\left(\eta^{2}-2\eta\epsilon+1\right)^{2}}<0,$ (59)
as $|\eta|<1$ and $|\epsilon|<1$. Therefore, $\widetilde{d}_{1}$ is the maximal point and $\widetilde{d}_{2}$ is the minimum point.
This is why $\widetilde{d}=0$ gives the maximal value in Fig. 3b. And in that figure, $\eta=0.005$, which leads to the minimal point $\widetilde{d}\approx 0.2$, so $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ decreases before $\widetilde{d}=0.2$ and increases after that point. The maximal and the minimal values of $\mathcal{H}_{d}$ can be obtained as
$\left.\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{1}}=$ $\frac{1-\epsilon^{2}}{\eta^{2}-2\eta\epsilon+1}\mathcal{H}_{d}^{({\rm opt})},$ (60)
$\left.\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{2}}=$ $\frac{\left(1-\epsilon^{2}\right)\left[\eta^{2}+w_{0}\left(\eta^{2}-2\eta\epsilon+1\right)+1\right]\mathcal{H}_{d}^{({\rm opt})}}{\left(1-\eta^{2}\right)^{2}\left(1-\epsilon^{2}\right)w_{0}+\left(1+\eta^{2}\right)\left(\eta^{2}-2\eta\epsilon+1\right)}.$
Interestingly, it can be verified that when $\widetilde{d}\rightarrow\pm\infty$, i.e., the separation between the two point sources is sufficiently large, the limit of $\mathcal{H}_{d}$ is
$\lim_{\widetilde{d}\rightarrow\pm\infty}=\frac{1-\epsilon^{2}}{\eta^{2}-2\eta\epsilon+1}\mathcal{H}_{d}^{({\rm opt})},$ (61)
which coincides with the maximal value of $\mathcal{H}_{d}$ at $\widetilde{d}=0$. This is why the value of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ cannot be larger than that at $\widetilde{d}=0$ when $\widetilde{d}$ goes beyond the minimal point in Fig. 3b though it increases with $\widetilde{d}$ in that regime.
When $\eta=0$, i.e., the widths of the two Gaussian point-spread functions are the same, it turns out that $\mathcal{H}_{d}$ has only one extreme point,
$\widetilde{d}_{3}=0.$ (62)
If we compute the second-order derivative of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$, it shows that
$\left.\partial_{\widetilde{d}}^{2}\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{3}}=\left(\frac{1}{\epsilon^{2}}-1\right)\mathcal{H}_{d}^{({\rm opt})}>0,$ (63)
as $|\epsilon|<1$. So, in sharp contrast to the case $\eta\neq 0$, the point $\widetilde{d}=0$ becomes the minimal point in the case $\eta=0$. This is why $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is the smallest at $\widetilde{d}=0$ in Fig. 2.
(a) $\epsilon=0$
(b) $\epsilon=0.25$
(c) $\epsilon=0.49$
(d) $\epsilon=0.64$
Figure 4: The ratio $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ with respect to the $\eta$ and $\widetilde{d}$ for different $\epsilon$. The plots show the discontinuity of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ at the point $\eta=0$, $\widetilde{d}=0$. Whenever $\epsilon\neq 0$, the limit of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is zero when $\widetilde{d}\rightarrow 0$ given $\eta=0$ and nonzero when $\eta\rightarrow 0$ given $\widetilde{d}=0$, so the limit $\widetilde{d}\rightarrow 0$ and the limit $\eta\rightarrow 0$ does not commute for the precision $\mathcal{H}_{d}$ in this case. When $\epsilon=0$, however, the values of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ in these two limits coincide, so the discontinuity of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ does not occur in this case.
Finally, it is worth noting that the precision $\mathcal{H}_{d}$ for the case $\sigma_{1}=\sigma_{2}$ and $\widetilde{d}\rightarrow 0$ in Fig. 2 behaves dramatically different from that for $\widetilde{d}=0$ and $\eta\rightarrow 0$ in Fig. 3a. The mechanism of this difference lies in the discontinuity of $\mathcal{H}_{d}$ at $\widetilde{d}=0$, $\eta=0$. It can be verified that
$\lim_{\widetilde{d}\rightarrow 0}\lim_{\eta\rightarrow 0}\mathcal{H}_{d}=$ $0,$ (64)
$\lim_{\eta\rightarrow 0}\lim_{\widetilde{d}\rightarrow 0}\mathcal{H}_{d}=$ $(1-\epsilon^{2})\mathcal{H}_{d}^{({\rm opt})},$
according to Eq. (47) and (51) whenever $\epsilon\neq 0$, so
$\lim_{\widetilde{d}\rightarrow 0}\lim_{\eta\rightarrow 0}\mathcal{H}_{d}\neq\lim_{\eta\rightarrow 0}\lim_{\widetilde{d}\rightarrow 0}\mathcal{H}_{d}.$ (65)
In Fig. 4, the precision $\mathcal{H}_{d}$ is plotted with respect to $\widetilde{d}$ and $\eta$ for different $\epsilon$ to illustrate this discontinuity. Hence, $\mathcal{H}_{d}$ has different behaviors with respect to $\epsilon$ in the above two cases.
VI Conclusions
The superresolution is an important quantum technique to surpass the Rayleigh’s criterion in resolving two closely located point sources, manifesting the power of quantum estimation theory in the field of quantum imaging. Nevertheless, the vanishing of the estimation precision when the photon numbers of the two point sources are unknown imposes limitations on its applicability. This work studies the condition for the superresolution technique to work by analyzing the competing roles of the difference in the photon numbers and the difference in the point-spread functions between the two optical sources as well as the separation itself in determining the estimation precision of the separation. The results not only provide conditions for the realizability of quantum superresolution, but also show various precision limits of superresolution in different regimes of parameters, which indicates that the competition between the difference of the photon numbers, the difference of the point-spread functions and the magnitude of the separation plays a key role in the validity of superresolution. We hope this work can provide new insights into the quantum theory of superresolution and extend the applications of quantum superresolution in practice.
Acknowledgements.
The authors acknowledge the helpful discussions with Jingyi Fan and Yutong Huang. This work is supported by the National Natural Science Foundation of China (Grant No. 12075323).
Appendix A Precision limit of direct imaging
Suppose the probability distribution of a single photon of position $x$ on the image plane is
$\Lambda(x)=\frac{1-\epsilon}{2}|\psi_{1}(x)|^{2}+\frac{1+\epsilon}{2}|\psi_{2}(x)|^{2}.$ (66)
Here, $\psi_{1}(x)$ and $\psi_{2}(x)$ are the point-spread functions of the two sources, and $\epsilon$ is the ratio between the difference and the total of the photon numbers of the two sources,
$\epsilon=\frac{N_{2}-N_{1}}{N_{2}+N_{1}}.$ (67)
In order to quantify the precision limit of estimating the separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which demonstrates that over all possible estimation strategies, the covariance matrix $C$ for the estimates of the parameters $g_{1},\cdots,g_{n}$ is lower bounded by
$C[g_{1},...,g_{n}]\geq M^{-1}\mathcal{J}^{-1}[g_{1},...,g_{n}].$ (68)
The $\geq$” sign represents semi-definite positivity of matrix, $M$ is the number of samples, and $\mathcal{J}[g_{1},...,g_{n}]$ denotes the Fisher information matrix defined as
[…]
where $P(x)$ is a probability distribution of random variable $x$ dependent on the parameters $g_{1},\cdots,g_{n}$. The benefit of the Cramér-Rao inequality is that its lower bound can always be asymptotically saturated by the maximum likelihood estimation strategy when $M$ is sufficiently large [15].
For the task of resolving two point sources, the positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate, and $P(x)$ is the position distribution $\Lambda(x)$ (66) of the photon from the incoherent sources.
For the direct imaging with an ideal continuum photodetector in the image plane, if the intensities and the point-spread functions of the two optical sources are equal, the position distribution (66) of the arrival single photon can be rewritten as
$\Lambda(x)=\frac{1}{2}I(x-\bar{X}+d/2)+\frac{1}{2}I(x-\bar{X}-d/2),$ (70)
where $I(x)$ is defined as
$I(x)=|\psi(x)|^{2}.$ (71)
When the separation $d$ is small, one can expand $\Lambda(x)$ to the second order of $d$,
$\Lambda(x)=I(x-\bar{X})+\frac{1}{8}d^{2}I^{\prime\prime}(x-\bar{X}).$ (72)
Then according Eq.(69), the classical Fisher information matrix for the parameters $\bar{X}$, $d$ can be obtained as
$\mathcal{J}=\begin{bmatrix}\alpha_{1}-\frac{\alpha_{2}d^{2}}{8}&0\\ 0&\frac{\alpha_{2}d^{2}}{16}\end{bmatrix},$ (73)
where
[…]
If one is interested in the parameter $d$ only (though $\bar{X}$ is still unknown), a weight matrix $W={\rm Diag}\{0,1\}$ can be multiplied to the both sides of the Cramér-Rao inequality (68), and the variance of the estimate of $d$ is lower bounded by
${\rm Var}(\hat{d})$ $\geqslant{\rm Tr}(W\mathcal{J}^{-1}),$ (75)
where $\hat{d}$ denotes the estimate of the separation $d$. Considering the variance may diverge when $d\rightarrow 0$, we use the reciprocal of the variance to quantify the estimation precision of $d$. Thus the estimation precision of the separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{8}d^{2}\alpha_{2}\text{{\hbox{N_{}}}}.$_ . (76)
Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the separation by direct imaging when the two point sources are sufficiently close to each other. This is called the Rayleigh’s curse [20] for direct imaging of two identical optical point sources.
The above result can be generalized to the case that the two point-spread functions are the same but the intensities are unequal. In this case, the position distribution of single photon can be written as
$\Lambda(x)=\frac{1-\epsilon}{2}I(x-\bar{X}+d/2)+\frac{1+\epsilon}{2}I(x-\bar{X}-d/2).$ (77)
Expanding the $\Lambda(x)$ to the second order of $d$, one can obtain
$\Lambda(x)$ $=I(x-\bar{X})-\frac{1}{2}d\epsilon I^{\prime}(x-\bar{X})+\frac{1}{8}d^{2}I^{\prime\prime}(x-\bar{X}).$ (78)
The Fisher information matrix $\mathcal{J}$ with respect to the parameters $\bar{X}$, $d$ can be obtained, and the entries are
$\mathcal{J}=\begin{bmatrix}\alpha_{1}+\frac{1}{8}d^{2}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]&\frac{\alpha_{1}\epsilon}{2}+\frac{\epsilon d^{2}}{16}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]\\ \frac{\epsilon\alpha_{1}}{2}+\frac{\epsilon d^{2}}{16}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]&\frac{\alpha_{1}\epsilon^{2}}{4}+\frac{d^{2}}{32}\left[\alpha_{2}\left(2-5\epsilon^{2}\right)+2\beta\epsilon^{4}\right]\end{bmatrix},$ (79)
[…]
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{N_{{\rm tot}}\alpha_{2}}{16}d^{2}\left(1-\epsilon^{2}\right)^{2},$ (80)
which means that no information about the separation $d$ can be acquired by direct imaging when $d\rightarrow 0$ in this generalized case.
Appendix B Multiparameter quantum metrology theory for superresolution
The quantum parameter estimation theory considers the estimation precision for unknown parameters in a parameter-dependent quantum state by quantum measurements. The core of the quantum parameter estimation theory is the quantum Cramér-Rao inequality. For the single-parameter case, the quantum Cramér-Rao inequality tells that the variance of the estimate of the unknown parameter is lower bounded by the inverse of the quantum Fisher information which maximizes the classical Fisher information over all possible quantum measurements on the parameter-dependent quantum state. For the multi-parameter case, the variance and the quantum Fisher information are generalized to the covariance matrix and the quantum Fisher information matrix respectively between the different parameters.
In detail, if one measures a quantum system with its density matrix $\hat{\rho}$ dependent on unknown parameters $g_{1},\cdots,g_{n}$ to determine these parameters, the covariance matrix for the estimates of these parameters is bounded by the following quantum Cramér-Rao’s inequality,
$C[g_{1},\cdots,g_{n}]\geq M^{-1}\mathcal{Q}^{-1}[g_{1},\cdots,g_{n}],$ (81)
where $M$ is the number of the quantum states, $C[g_{1},\cdots,g_{n}]$ is the covariance matrix for the estimates of $g_{1},\cdots,g_{n}$, $\mathcal{Q}[g_{1},\cdots,g_{n}]$ is the quantum Fisher information matrix about $g_{1},\cdots,g_{n}$ with entries
$\mathcal{Q}_{ij}(\hat{\rho})=\frac{1}{2}{\rm Tr}[\hat{\rho}\{\hat{\mathcal{L}}_{i},\hat{\mathcal{L}}_{j}\}],$ (82)
and ^$\hat{\mathcal{L}}_{i}$, ^$\hat{\mathcal{L}}_{j}$ are the symmetric logarithmic derivatives (SLD) of the density operator $\hat{\rho}$,
$\frac{1}{2}(\hat{\mathcal{L}}_{i}\hat{\rho}+\hat{\rho}\hat{\mathcal{L}}_{i})=\partial_{g_{i}}\hat{\rho}.$ (83)
To obtain the quantum Fisher information matrix for $\hat{\rho}$ which depends on the unknown parameters $g_{i}$, one can decompose the density matrix $\hat{\rho}$ as
$\hat{\rho}=\sum_{k}\lambda_{k}|\lambda_{k}\rangle\langle\lambda_{k}|,$ (84)
where $\lambda_{k}$ and $|\lambda_{k}\rangle$ are eigenvalue and eigenstate of $\hat{\rho}$. The matrix elements of both sides of Eq. (83) with respect to the eigenbasis $|\lambda_{k}\rangle$ can be obtained as
$\langle\lambda_{k}|\partial_{g_{i}}\hat{\rho}|\lambda_{h}\rangle=\frac{1}{2}\left(\langle\lambda_{k}|\hat{\rho}\hat{\mathcal{L}}_{i}|\lambda_{h}\rangle+\langle\lambda_{k}|\hat{\mathcal{L}}_{i}\hat{\rho}|\lambda_{h}\rangle\right),$ (85)
which leads to
^$\hat{\mathcal{L}}_{i}=\sum_{\lambda_{k}+\lambda_{h}\neq 0}\frac{2\langle\lambda_{k}|\partial_{g_{i}}\hat{\rho}|\lambda_{h}\rangle}{\lambda_{k}+\lambda_{h}}|\lambda_{k}\rangle\langle\lambda_{h}|.$ (86)
Thus, the elements of the quantum Fisher information matrix can be obtained as
$\mathcal{Q}_{ij}(\hat{\rho})=2\sum_{\lambda_{k}+\lambda_{l}\neq 0}\frac{\langle\lambda_{k}|\partial_{g_{i}}\hat{\rho}|\lambda_{l}\rangle\langle\lambda_{l}|\partial_{g_{j}}\hat{\rho}|\lambda_{k}\rangle}{\lambda_{k}+\lambda_{l}}.$ (87)
As the density matrix (1) of a single photon is low-rank, it is more convenient to compute the quantum Fisher information matrix on the support of $\hat{\rho}$,
$\mathcal{Q}_{ij}(\hat{\rho})=$ $\sum_{\lambda_{k}\neq 0}\frac{4\langle\lambda_{k}|\partial_{g_{i}}\hat{\rho}\partial_{g_{j}}\hat{\rho}|\lambda_{k}\rangle}{\lambda_{k}}$ (88)
$+\sum_{\lambda_{k},\lambda_{h}\neq 0}2\left(\frac{1}{\lambda_{k}+\lambda_{h}}-\frac{1}{\lambda_{k}}-\frac{1}{\lambda_{h}}\right)$
×$\times\langle\lambda_{h}|\partial_{g_{i}}\hat{\rho}|\lambda_{k}\rangle\langle\lambda_{k}|\partial_{g_{j}}\hat{\rho}|\lambda_{h}\rangle,$
which is a variant of the result in Ref. [109].
For the superresolution of two point sources, the positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$, are the unknown parameters to estimate, so the multi-parameter quantum Cramér-Rao bound (81) tells that
$C[\bar{X},\,d]\geq N_{{\rm tot}}^{-1}\mathcal{Q}^{-1}[\bar{X},\,d],$ (89)
where $N_{{\rm tot}}$ is the total number of photons $N_{1}+N_{2}$, $C[\bar{X},\,d]$ is the covariance matrix for the estimates of $\bar{X},\,d$, $\mathcal{Q}[\bar{X},\,d]$ is the quantum Fisher information matrix about $\bar{X},\,d$ with entries in Eq. (82).
If one is interested in only one or a few unknown parameters, e.g., the separation as in the superresolution, a positive semidefinite weight matrix $W$ can be imposed on both sides of the quantum Cramér-Rao inequality (89), then the matrix quantum Cramér-Rao bound turns to a scalar form,
${\rm Tr}(WC)\geq N_{{\rm tot}}^{-1}{\rm Tr}(W\mathcal{Q}^{-1}).$ (90)
Note that for quantum multiparameter estimation, the optimal measurements for different parameters are not always compatible, so the matrix quantum Cramér-Rao bound is not always achievable. The saturation condition given by Ragy et al. [100] for the above scalar quantum Cramér-Rao bound is
${\rm Tr}(\hat{\rho}[\hat{\mathcal{L}}_{i},\hat{\mathcal{L}}_{j}])={\rm Im[Tr}(\hat{\rho}\hat{\mathcal{L}}_{i}\hat{\mathcal{L}}_{j})]=0,$ (91)
as the scalar quantum Cramér-Rao bound $N_{tot}^{-1}{\rm Tr}(W\mathcal{Q}^{-1})$ coincides with the Holevo bound in this case and the latter is always asymptotically achievable when $N_{{\rm tot}}\rightarrow\infty$. As the point-spread functions $\psi_{i}(x)$ are real in the superresolution problem, the density matrix $\hat{\rho}$ and the eigenvalues and eigenstates of $\hat{\rho}$ are also real in this case as will be shown in the following sections, so the compatibility condition (91) is saturated, and the scalar quantum Cramér-Rao bound (90) can be achieved by the maximum-likelihood estimation strategy for $N_{{\rm tot}}\rightarrow\infty$.
Appendix C Precision for superresolution of two identical optical sources
For two incoherent optical sources if one photon is emitted, the quantum density operator for the single photon of the optical field on the image plane can be written as
$\hat{\rho}=\frac{1-\epsilon}{2}{\rm e}^{-iX_{1}\hat{P}}|\psi\rangle\langle\psi|{\rm e}^{iX_{1}\hat{P}}+\frac{1+\epsilon}{2}{\rm e}^{-iX_{2}\hat{P}}|\psi\rangle\langle\psi|{\rm e}^{iX_{2}\hat{P}},$ (92)
where we define $\hat{P}=-i\partial_{x}$ for the sake of convenience. The density matrix $\hat{\rho}$ can be infinite-dimensional in principle, but has only two nonzero eigenvalues which can be obtained as
$\lambda_{1,2}=\frac{1}{2}\left(1\mp\sqrt{\epsilon^{2}+\delta^{2}\left(1-\epsilon^{2}\right)}\right).$ (93)
The associated eigenstates of $\hat{\rho}$ are
$|\lambda_{1,2}\rangle$ $=\frac{1}{\sqrt{\mathscr{Q}_{1,2}}}\times\{{\rm e}^{-iX_{1}\hat{P}}|\psi\rangle\mp\frac{\left(\sqrt{\epsilon^{2}+\delta^{2}\left(1-\epsilon^{2}\right)}\mp\epsilon\right)}{\delta(1-\epsilon)}{\rm e}^{-iX_{2}\hat{P}}|\psi\rangle\},$ (94)
where $\mathscr{Q}_{1,2}$ are the normalization constants,
$\mathscr{Q}_{1,2}$ $=\left(1-\delta^{2}\right)\left(1+\frac{\delta^{2}\left(1-\delta^{2}\right)(1-\epsilon)^{2}}{\left(\sqrt{\delta^{2}+\epsilon^{2}-\delta^{2}\epsilon^{2}}\pm\delta^{2}(1-\epsilon)\pm\epsilon\right)^{2}}\right),$ (95)
and because point-spread functions are assumed to be real, so it can simplify to
$\delta=\langle\psi|{\rm e}^{-id\hat{P}}|\psi\rangle=\langle\cos(d\hat{P})\rangle.$ (96)
C.1 With known arbitrary photon numbers
When the photon numbers $N_{1}$ and $N_{2}$ of the two point sources are known, one can derive the quantum Fisher information matrix $\mathcal{Q}$ for parameters $\bar{X},\,d$ by plugging the above eigenvalues and eigenstates of $\hat{\rho}$ into Eq. (88),
$\mathcal{Q}=\begin{bmatrix}4\kappa-4\gamma^{2}\left(1-\epsilon^{2}\right)&2\kappa\epsilon\\ 2\kappa\epsilon&\kappa\end{bmatrix}.$ (97)
Taking the inverse of the quantum Fisher information matrix (105), the estimation precision of the separation $d$ can be worked out as
$\mathcal{H}_{d}=N_{{\rm tot}}/(\mathcal{Q}^{-1})_{22}=\frac{N_{{\rm tot}}\kappa\left(1-\epsilon^{2}\right)\left(\kappa-\gamma^{2}\right)}{\kappa-\gamma^{2}\left(1-\epsilon^{2}\right)},$ (98)
where
$\kappa=\langle\psi|\hat{P}^{2}|\psi\rangle=\langle\hat{P}{}^{2}\rangle,\;\gamma=i\langle\psi|\hat{P}{\rm e}^{-id\hat{P}}|\psi\rangle=\langle\sin(d\hat{P})\hat{P}\rangle.$ (99)
When the separation goes to zero, the estimation separation precision $\mathcal{H}_{d}$ becomes
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=N_{{\rm tot}}\left(1-\epsilon^{2}\right)\langle\hat{P}^{2}\rangle.$ (100)
For two Gaussian point-spread functions, the estimation precision for the separation $d$ is
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left(4\sigma^{2}{\rm e}^{\frac{d^{2}}{4\sigma^{2}}}-d^{2}\right)}{4\sigma^{2}\left(4\sigma^{2}{\rm e}^{\frac{d^{2}}{4\sigma^{2}}}-d^{2}\left(1-\epsilon^{2}\right)\right)}.$ (101)
In the limit $d\rightarrow 0$, the estimation precision of the separation (100) can be simplified to
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)}{4\sigma^{2}},$ (102)
which shows $\mathcal{H}_{d}$ keeps nonzero when $d$ is zero.
When the photon numbers of the two optical sources are the same, $\epsilon=0$, $N_{1}=N_{2}=N$ and $N_{{\rm tot}}=2N$, the general result (100) for the estimation precision of $d$ in the limit $d\rightarrow 0$ can be simplified to
$\mathcal{H}_{d}=2N\langle\hat{P}^{2}\rangle.$ (103)
And for the Gaussian case, the result becomes
$\mathcal{H}_{d}=\frac{N}{2\sigma^{2}}.$ (104)
This recovers the results for the original superresolution scheme in the paper by Tsang et al. [20].
C.2 Unknown arbitrary photon numbers
For two incoherent optical sources with unknown photon numbers, according to Eq. (88), one can derive the quantum Fisher information matrix $\mathcal{Q}$ for parameters $\bar{X},\,d,\,\epsilon$ as
$\mathcal{Q}=\left[\begin{array}[]{ccc}4\kappa-4\gamma^{2}\left(1-\epsilon^{2}\right)&2\kappa\epsilon&2\gamma\delta\\ 2\kappa\epsilon&\kappa&0\\ 2\gamma\delta&0&\frac{1-\delta^{2}}{1-\epsilon^{2}}\end{array}\right],$ (105)
where $\kappa$, $\gamma$ are defined in Eq. (99) and $\delta$ is defined in Eq. (96).Taking the inverse of the quantum Fisher information matrix $\mathcal{Q}$ (105), the estimation precision of the separation $d$ can be worked out as
$\mathcal{H}_{d}=N_{{\rm tot}}/(\mathcal{Q}^{-1})_{22}=\frac{N_{{\rm tot}}\kappa\left(1-\epsilon^{2}\right)\left[\left(1-\delta^{2}\right)\kappa-\gamma^{2}\right]}{(1-\delta^{2})\kappa-\gamma^{2}\left(1-\epsilon^{2}\right)}.$ (106)
If the two photon numbers $N_{1}$, $N_{2}$ are different and $N_{2}-N_{1}$ keeps finite, the precision of the separation (106) can be simplified to
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right){\rm Var}(\hat{P}^{2})d^{2}}{4\epsilon^{2}},$ (107)
which is of order $O(d^{2})$, where ${\rm Var}(\hat{P}^{2})=\langle\hat{P}^{4}\rangle-\langle\hat{P}{}^{2}\rangle^{2}$. This indicates that estimation precision for the separation vanishes when $d$ is small, in drastic constraint to the above results for resolving two identical incoherent optical sources with known photon numbers. So the superresolution scheme does not work and the Rayleigh’s curse takes effect again in this case, which is sometimes called the resurgence of Rayleigh’s curse. In order to study how to address this issue, we need to take a deeper insight into the estimation precision $\mathcal{H}_{d}$ by expanding the numerator and denominator of $\mathcal{H}_{d}$ (106) to the fourth order of $d$,
$\mathcal{H}_{d}=\frac{A_{0}+A_{2}d^{2}+A_{4}d^{4}}{B_{0}+B_{2}d^{2}+B_{4}d^{4}},$ (108)
with the expansion coefficients
$A_{0}$ $=0,\;B_{0}=0,$ (109)
$A_{2}$ $=0,\;B_{2}=12\epsilon^{2}\langle\hat{P}{}^{2}\rangle,$
$A_{4}$ $=3N_{{\rm tot}}(1-\epsilon^{2})\langle\hat{P}{}^{2}\rangle{\rm Var}(\hat{P}^{2}),$
$B_{4}$ $=3{\rm Var}(\hat{P}^{2})-4\epsilon^{2}\langle\hat{P}^{4}\rangle,$
which turns out to be
$\mathcal{H}_{d}=\frac{3N_{{\rm tot}}d^{2}\left(1-\epsilon^{2}\right)\langle\hat{P}{}^{2}\rangle{\rm Var}(\hat{P}^{2})}{12\epsilon^{2}\langle\hat{P}{}^{2}\rangle+d^{2}\left[3{\rm Var}(\hat{P}^{2})-4\epsilon^{2}\langle\hat{P}^{4}\rangle\right]}.$ (110)
Obviously, when the photon numbers of the two incoherent point sources are the same, i.e., $\epsilon=0$, the leading constant term of the denominator of $\mathcal{H}_{d}$ vanishes and both the numerator and the denominator of $\mathcal{H}_{d}$ are of order $O(d^{2})$, so $\mathcal{H}_{d}$ keeps nonzero and independent of $d$ when $d\rightarrow 0$, which is the quantum superresolution of two incoherent point sources. In contrast, if the photon numbers of the two optical sources are different, i.e., $N_{1}\neq N_{2}$, the magnitude of the numerator of $\mathcal{H}_{d}$ is $O(d^{2})$ but the magnitude of the denominator of $\mathcal{H}_{d}$ is $O(1)$ if the difference between the two photon numbers keeps finite. In this case when $d\rightarrow 0$ the separation precision vanishes, which is the limitation imposed by the Rayleigh’s curse.
Take the Gaussian point-spread functions as an example. The estimation precision $\mathcal{H}_{d}$ (106) becomes
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left[d^{2}+4\sigma^{2}\left(1-{\rm e}^{\frac{d^{2}}{4\sigma^{2}}}\right)\right]}{4\sigma^{2}\left[4\sigma^{2}\left(1-{\rm e}^{\frac{d^{2}}{4\sigma^{2}}}\right)+d^{2}\left(1-\epsilon^{2}\right)\right]}.$ (111)
If we further define a dimensionless separation between the two Gaussian point sources as
$\widetilde{d}=\frac{d}{\sigma},$ (112)
the estimation precision $\mathcal{H}_{d}$ (111) can be simplified to
$\mathcal{H}_{d}=\frac{\left(4e^{\frac{\widetilde{d}^{2}}{4}}-\widetilde{d}^{2}-4\right)\left(1-\epsilon^{2}\right)}{4e^{\frac{\widetilde{d}^{2}}{4}}-\widetilde{d}^{2}\left(1-\epsilon^{2}\right)-4}\mathcal{H}_{d}^{({\rm opt})},$ (113)
where $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision of $d$ when the photon numbers of the two sources are the same, i.e., $N_{1}=N_{2}=N$,
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N}{2\sigma^{2}}.$ (114)
This implies that the ratio between the estimation precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ is determined by the parameters $\epsilon$ and $r$ only.
By expanding both the numerator and denominator of Eq. (111) to the fourth order of $d$, one can have
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)d^{2}}{32\sigma^{4}\epsilon^{2}+4d^{2}\sigma^{2}},$ (115)
where a common $d^{2}$ has been reduced from both the numerator and denominator. It shows $\mathcal{H}_{d}$ keeps nonzero when $d$ is small if $\epsilon=0$ which is the result of quantum superresolution and vanishes if $\epsilon$ is nonzero and keeps finite which is the result of Rayleigh’s criterion.
The above analysis givens an important implication that the magnitude of $\epsilon$ plays a critical role in determining the effect of superresolution. If $N_{1}$ and $N_{2}$ are different but the magnitude of $\epsilon$ is of order $O(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle})$ or higher, the order of the denominator of $\mathcal{H}_{d}$ is still $O(d^{2})$ then and $\mathcal{H}_{d}$ will not vanish in this case. This immediately leads to the condition for the superresolution with two identical incoherent optical sources in the main text.
Appendix D Precision for resolving two arbitrary point sources
D.1 General point-spread functions
For two arbitrary optical point sources, the density operator of a single photon can be written as
$\hat{\rho}=\frac{1-\epsilon}{2}{\rm e}^{-iX_{1}\hat{P}}|\psi_{1}\rangle\langle\psi_{1}|{\rm e}^{iX_{1}\hat{P}}+\frac{1+\epsilon}{2}{\rm e}^{-iX_{2}\hat{P}}|\psi_{2}\rangle\langle\psi_{2}|{\rm e}^{iX_{2}\hat{P}}.$ (116)
To obtain the quantum Fisher information matrix, we need to work out the eigenvalues and eigenstates of $\hat{\rho}$. Note that $\hat{\rho}$ can be infinite-dimensional in principle, but has only two nonzero eigenvalues, which turn out to be
$\lambda_{1,2}=\frac{1}{2}\left(1\mp\sqrt{\epsilon^{2}+\delta^{2}\left(1-\epsilon^{2}\right)}\right),$ (117)
and the associated eigenstates $|\lambda_{1,2}\rangle$ are
$|\lambda_{1,2}\rangle=\frac{1}{\sqrt{\mathscr{Q}_{1,2}}}\times\{{\rm e}^{-iX_{1}\hat{P}}|\psi_{1}\rangle\mp\frac{\left(\sqrt{\epsilon^{2}+\delta^{2}\left(1-\epsilon^{2}\right)}\mp\epsilon\right)}{\delta(1-\epsilon)}{\rm e}^{-iX_{2}\hat{P}}|\psi_{2}\rangle\},$ (118)
where $\delta$ is the overlap between the point-spread functions of the two sources,
$\delta=\langle\psi_{1}|{\rm e}^{iX_{1}\hat{P}}{\rm e}^{-iX_{2}\hat{P}}|\psi_{2}\rangle=\langle{\rm e}^{-id\hat{P}}\rangle_{12}=\langle\cos(d\hat{P})\rangle_{12},$ (119)
with $\langle\cdot\rangle_{i}=\langle\psi_{i}|\cdot|\psi_{i}\rangle$ and $\langle\cdot\rangle_{12}=\langle\psi_{1}|\cdot|\psi_{2}\rangle$, $\mathscr{Q}_{1,2}$ are the normalization constants,
$\mathscr{Q}_{1,2}=\left(1-\delta^{2}\right)\left(1+\frac{\delta^{2}\left(1-\delta^{2}\right)(1-\epsilon)^{2}}{\left(\sqrt{\delta^{2}+\epsilon^{2}-\delta^{2}\epsilon^{2}}\pm\delta^{2}(1-\epsilon)\pm\epsilon\right)^{2}}\right).$ (120)
When the photon numbers of the two optical sources are unknown, one needs to estimate the ratio $\epsilon$ of the difference between the two photon numbers to the total photon number in addition to the centroid $\bar{X}$ and the separation $d$ of the two sources. The quantum Fisher information matrix $\mathcal{Q}$ for parameters $\bar{X},\,d,\,\epsilon$ can be derived by Eq. (88) as
$\mathcal{Q}=\left[\begin{array}[]{ccc}2\kappa_{{\rm tot}}(1+\chi\epsilon)-4\gamma^{2}\left(1-\epsilon^{2}\right)&\kappa_{{\rm tot}}(\chi+\epsilon)&2\gamma\delta\\ \kappa_{{\rm tot}}(\chi+\epsilon)&\frac{1}{2}\kappa_{{\rm tot}}(1+\chi\epsilon)&0\\ 2\gamma\delta&0&\frac{1-\delta^{2}}{1-\epsilon^{2}}\end{array}\right],$ (121)
where
$\kappa_{i}=$ $\langle\partial_{X_{i}}\psi_{i}(x-X_{i})|\partial_{X_{i}}\psi_{i}(x-X_{i})\rangle=\langle\psi_{i}|P^{2}|\psi_{i}\rangle=\langle\hat{P}{}^{2}\rangle_{i},\;i=1,2,$ (122)
$\gamma=$ $\langle\partial_{X_{1}}\psi_{1}(x-X_{1})|\psi_{2}(x-X_{2})\rangle=i\langle\psi_{1}|P{\rm e}^{-idP}|\psi_{2}\rangle=\langle\sin(d\hat{P})\hat{P}\rangle_{12},$
$\kappa_{{\rm tot}}=$ $\langle\hat{P}{}^{2}\rangle_{1}+\langle\hat{P}{}^{2}\rangle_{2},$
$\chi=$ $\frac{\langle\hat{P}{}^{2}\rangle_{2}-\langle\hat{P}{}^{2}\rangle_{1}}{\langle\hat{P}{}^{2}\rangle_{1}+\langle\hat{P}{}^{2}\rangle_{2}}.$
By taking the inverse of the matrix $\mathcal{Q}$, the quantum Cramér-Rao’s bound for estimation precision of the separation $d$ can be derived as
$\mathcal{H}_{d}=N_{{\rm tot}}/(Q^{-1})_{22}=\frac{N_{{\rm tot}}\kappa_{{\rm tot}}\left(1-\epsilon^{2}\right)\left[\left(1-\delta^{2}\right)\left(1-\chi^{2}\right)\kappa_{{\rm tot}}-2\gamma^{2}(1+\chi\epsilon)\right]}{2\kappa_{{\rm tot}}\left(1-\delta^{2}\right)(\chi\epsilon+1)-4\gamma^{2}\left(1-\epsilon^{2}\right)}.$ (123)
Now, one can expand $\delta$ and $\gamma$ to the fourth order of $d$ (note that $\kappa_{1}$, $\kappa_{2}$ are independent of $d$ as can be seen from Eq. (122)),
$\delta$ $=\langle\psi_{1}|\psi_{2}\rangle-\frac{1}{2}\langle\hat{P}{}^{2}\rangle_{12}d^{2}+\frac{1}{24}\langle\hat{P}^{4}\rangle_{12}d^{4},$ (124)
$\gamma$ $=d\langle\hat{P}{}^{2}\rangle_{12}-\frac{1}{6}d^{3}\langle\hat{P}^{4}\rangle_{12},$
Plugging Eq. (124) into $\mathcal{H}_{d}$ (123), the numerator and the denominator of $\mathcal{H}_{d}$ can be expanded to the fourth order of $d$ respectively, which gives
$\mathcal{H}_{d}=\frac{A_{0}+A_{2}d^{2}+A_{4}d^{4}}{B_{0}+B_{2}d^{2}+B_{4}d^{4}},$ (125)
where
$A_{0}$ $=-12\left(1-\chi^{2}\right)\left(1-\epsilon^{2}\right)\left(1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}\right)N_{{\rm tot}}\kappa_{{\rm tot}},$ (126)
$B_{0}$ $=-24(1+\chi\epsilon)\left(1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}\right),$
$A_{2}$ $=12N_{{\rm tot}}\kappa_{{\rm tot}}^{2}\xi\left(1-\epsilon^{2}\right)\left[2\xi-\langle\psi_{1}|\psi_{2}\rangle+\chi^{2}\langle\psi_{1}|\psi_{2}\rangle+2\xi\chi\epsilon\right],$
$B_{2}$ $=24\xi\kappa_{{\rm tot}}\left[\left(2\xi-\langle\psi_{1}|\psi_{2}\rangle\right)-\chi\epsilon\langle\psi_{1}|\psi_{2}\rangle-2\xi\epsilon^{2}\right],$
$A_{4}$ $=\left(1-\epsilon^{2}\right)N_{{\rm tot}}\kappa_{{\rm tot}}\left[3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}-8\xi\chi\epsilon\langle\hat{P}^{4}\rangle_{12}-\chi^{2}\left(\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+3\xi^{2}\kappa_{{\rm tot}}^{2}\right)\right],$
$B_{4}$ $=2\left[3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+8\xi\epsilon^{2}\langle\hat{P}^{4}\rangle_{12}+\chi\epsilon\left(\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+3\xi^{2}\kappa_{{\rm tot}}^{2}\right)\right],$
and
$\xi=\frac{\langle\hat{P}{}^{2}\rangle_{12}}{\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}}.$ (127)
If $\langle\psi_{1}|\psi_{2}\rangle\neq 1$ and $1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}$ keeps nonzero and finite, when $d\rightarrow 0$, $\mathcal{H}_{d}$ can be simplified to
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{N_{{\rm tot}}\kappa_{{\rm tot}}\left(1-\chi^{2}\right)\left(1-\epsilon^{2}\right)}{2(1+\chi\epsilon)}.$ (128)
For this case, the Rayleigh’s criterion is violated and the superresolution works. Actually in this case, only $A_{0}$ and $B_{0}$ are involved.
When the difference between the two point-spread functions and the difference between the two photon numbers are comparable to the magnitude of the separation $d$, the higher-order terms in Eq. (125) will also contribute to $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ And the results are summarized in Table 1 of the main text.
D.2 Gaussian point-spread functions
For two Gaussian point-spread functions with positions $X_{1},X_{2}$ and widths $\sigma_{1},\sigma_{2}$, the parameters $\delta,\gamma,\kappa_{1},\kappa_{2}$ can be obtained as
$\delta$ $=\sqrt{\frac{2\sigma_{1}\sigma_{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}}\exp\left[-\frac{d^{2}}{4\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)}\right],$ (129)
$\gamma$ $=d\sqrt{\frac{\sigma_{1}\sigma_{2}}{2(\sigma_{1}^{2}+\sigma_{2}^{2})^{3}}}\exp\left[-\frac{d^{2}}{4\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)}\right],$
$\kappa_{1}$ $=\frac{1}{4\sigma_{1}^{2}},\;\kappa_{2}=\frac{1}{4\sigma_{2}^{2}},$
where $\sigma_{{\rm tot}}=\sigma_{1}+\sigma_{2}$. The estimation precision of the separation $d$ turns to be
$\mathcal{H}_{d}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left\{\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}\left[\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){\rm e}^{\frac{d^{2}}{2\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)}}-2\sigma_{2}\sigma_{1}\right]-d^{2}\left[\sigma_{1}^{3}\sigma_{2}(1+\epsilon)+\sigma_{1}\sigma_{2}^{3}(1-\epsilon)\right]\right\}}{2\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}\left[\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){\rm e}^{\frac{d^{2}}{2\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)}}-2\sigma_{1}\sigma_{2}\right]\left[\sigma_{1}^{2}(\epsilon+1)+\sigma_{2}^{2}(1-\epsilon)\right]-8\left(1-\epsilon^{2}\right)\sigma_{1}^{3}\sigma_{2}^{3}d^{2}}.$ (130)
For two Gaussian point-spread functions with different widths $\sigma_{1}$ and $\sigma_{2}$, the coefficients in Eq. (125) can be obtained as
$A_{0}=$ $8N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left(\sigma_{1}-\sigma_{2}\right){}^{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2},$ (131)
$B_{0}=$ $16\left(\sigma_{1}-\sigma_{2}\right){}^{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right],$
$A_{2}=$ $4N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left\{\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}-2\sigma_{2}\sigma_{1}\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right]\right\},$
$B_{2}=$ $8\left\{\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right){}^{2}\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right]-8\sigma_{1}^{3}\sigma_{2}^{3}\left(1-\epsilon^{2}\right)\right\},$
$A_{4}=$ $N_{{\rm tot}}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)\left(1-\epsilon^{2}\right),$
$B_{4}=$ $2\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right].$
When $d\rightarrow 0$, it can be simplified to
$\left.\mathcal{H}_{d}\right|_{d\rightarrow 0}=\frac{N_{{\rm tot}}\left(1-\epsilon^{2}\right)}{2\left[\sigma_{1}^{2}(1+\epsilon)+\sigma_{2}^{2}(1-\epsilon)\right]},$ (132)
which is Eq. (51) in the main text. The results of the estimation precision $\mathcal{H}_{d}$ in different regimes of the photon numbers and the widths of the two point sources are summarized in Table 2 of the main text.
If we denote
$\eta=\frac{\sigma_{2}-\sigma_{1}}{\sigma_{1}+\sigma_{2}},\quad\widetilde{d}=\frac{d}{(\sigma_{1}+\sigma_{2})/2},$ (133)
where $\widetilde{d}$ is the dimensionless separation extended from the case of two identical point-spread functions (112), the estimation precision $\mathcal{H}_{d}$ (130) can be simplified as
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)\left\{4\left(\eta^{2}+1\right)^{3}e^{\frac{\widetilde{d}^{2}}{4\eta^{2}+4}}-\left(1-\eta^{2}\right)\left[4\left(\eta^{2}+1\right)^{2}+\widetilde{d}^{2}\left(\eta^{2}-2\eta\epsilon+1\right)\right]\right\}}{4\left(\eta^{2}+1\right)^{2}\left(\eta^{2}-2\eta\epsilon+1\right)\left[\left(\eta^{2}+1\right)e^{\frac{\widetilde{d}^{2}}{4\eta^{2}+4}}-\left(1-\eta^{2}\right)\right]-\widetilde{d}^{2}\left(1-\eta^{2}\right)^{3}\left(1-\epsilon^{2}\right)}\mathcal{H}_{d}^{({\rm opt})},$ (134)
and $\mathcal{H}_{d}^{({\rm opt})}$ is the optimal estimation precision of $d$ when the photon numbers and the widths of the two Gaussian sources are the same, i.e., $N_{1}=N_{2}=N$ and $\sigma_{1}=\sigma_{2}=\sigma$,
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N}{2\sigma^{2}}.$ (135)
This implies that the ratio between the precision $\mathcal{H}_{d}$ and the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ is determined by the parameters $\epsilon$, $\eta$ and $\widetilde{d}$ only.
Appendix E Competing quantities in the validity of superresolution
E.1 Critical quantities
It can be seen from the results for two identical point sources that the validity of superresolution is highly sensitive to the magnitude of the leading term in the denominator of $\mathcal{H}_{d}$ that may be comparable to $d$ and it turns out that the ratio of the difference between the photon numbers of the two optical sources to the total photon number,
$\epsilon=\frac{N_{2}-N_{1}}{N_{2}+N_{1}}.$ (136)
determines the magnitude of that term which leads to quantum superresolution if it is of order $O(\widetilde{d})$ or even equal to zero. So, $\epsilon$ is a critical quantity that determines the validity of superresolution for two identical point-spread functions.
When the two point-spread functions differ slightly from each other, the difference between the two point-spread functions may change the magnitude of the leading terms in the numerator and denominator of $\mathcal{H}_{d}$ and render the additional terms in $\mathcal{H}_{d}$ (125) comparable to $\widetilde{d}$ which will compete with each other as well as with the ratio $\epsilon$ between the photon number difference and the total photon number in determining the validity of superresolution, so the difference of the point-spread functions can have significant effect on $\mathcal{H}_{d}$ and we need to investigate this effect in detail.
As a first step to analyze the effect of the difference between the point-spread functions, we figure out what quantities in Eq. (125) are determined by the differences between the two point-spread functions in the following.
For the zeroth-order terms $A_{0}$ and $B_{0}$ of the numerator and denominator of $\mathcal{H}_{d}$ (125), Eq. (126) tells that
$A_{0}$ $=-12N_{{\rm tot}}\kappa_{{\rm tot}}\left(1-\chi^{2}\right)\left(1-\epsilon^{2}\right)\left(\underline{1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}}\right),$ (137)
$B_{0}$ $=-24(1+\chi\epsilon)\left(\underline{1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}}\right),$
where $N_{{\rm tot}}$ is the total photon number,
$N_{{\rm tot}}=N_{1}+N_{2}.$ (138)
It can be seen that $A_{0}$ and $B_{0}$ vanish when the two point-spread functions are close to each other, i.e., $\psi_{1}(x)-\psi_{2}(x)\rightarrow 0$ or equivalently $\langle\psi_{1}|\psi_{2}\rangle\rightarrow 1$, so $1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}$ is also critical to the estimation precision $\mathcal{H}_{d}$. As $1-\left|\langle\psi_{1}|\psi_{2}\rangle\right|^{2}=\left(1-\langle\psi_{1}|\psi_{2}\rangle\right)\left(1+\langle\psi_{1}|\psi_{2}\rangle\right)$ and $1+\langle\psi_{1}|\psi_{2}\rangle\neq 0$ when $\psi_{1}(x)$ and $\psi_{2}(x)$ are close to each other, so for simplicity we choose $1-\langle\psi_{1}|\psi_{2}\rangle$ to be the quantity that is determined by the difference between the two point-spread functions and critical to the estimation precision $\mathcal{H}_{d}$.
For the second-order terms $A_{2}d^{2}$ and $B_{2}d^{2}$ in the numerator and denominator of $\mathcal{H}_{d}$ (125), the coefficients $A_{2}$ and $B_{2}$ can be rewritten as
$A_{2}=$ $12N_{{\rm tot}}\kappa_{{\rm tot}}^{2}\xi\left(1-\epsilon^{2}\right)\left[\underline{2\xi-\langle\psi_{1}|\psi_{2}\rangle}+\underline{\chi^{2}}\langle\psi_{1}|\psi_{2}\rangle+2\xi\underline{\chi}\epsilon\right],$ (139)
$B_{2}=$ $24\xi\kappa_{{\rm tot}}\left[\underline{2\xi-\langle\psi_{1}|\psi_{2}\rangle}-\underline{\chi}\epsilon\langle\psi_{1}|\psi_{2}\rangle-2\xi\epsilon^{2}\right],$
where
$\text{{\hbox{\kappa_{{\rm tot}}^{}}}}=\kappa_{1}+\kappa_{2}=\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}.$ (140)
When the two point-spread functions are close to each other, $\langle\psi_{1}|\psi_{2}\rangle\rightarrow 1$, $\chi\rightarrow 0$, $2\xi-\langle\psi_{1}|\psi_{2}\rangle=2\xi-1\rightarrow 0$ , which leads to $A_{2}\rightarrow 0$, $B_{2}\rightarrow 0$, so one can immediately find that $\chi$ and $\frac{1}{2}-\xi$ are the two additional quantities that are determined by the difference between the two point-spread functions and can change the estimation precision of the separation $d$.
For the fourth-order terms $A_{4}d^{4}$ and $B_{4}d^{4}$ in the numerator and denominator of $\mathcal{H}_{d}$ (125), the coefficients $A_{4}$ and $B_{4}$ can be rewritten as
$A_{4}=$ $\left(1-\epsilon^{2}\right)N_{{\rm tot}}\kappa_{{\rm tot}}\bigg{[}3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}-8\xi\chi\epsilon\langle\hat{P}^{4}\rangle_{12}-\chi^{2}\left(\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+3\xi^{2}\kappa_{{\rm tot}}^{2}\right)\bigg{]},$ (141)
$B_{4}=$ $2\left[3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+8\xi\epsilon^{2}\langle\hat{P}^{4}\rangle_{12}+\chi\epsilon\left(\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}+3\xi^{2}\kappa_{{\rm tot}}^{2}\right)\right].$
When the two point-spread functions are close to each other, $\langle\psi_{1}|\psi_{2}\rangle\rightarrow 1$ and $\chi\rightarrow 0$, $\xi\rightarrow\frac{1}{2}$, $\epsilon\rightarrow 0$, leading to
$A_{4}\rightarrow$ $N_{{\rm tot}}\kappa_{{\rm tot}}\left(3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}\right),$ (142)
$B_{4}\rightarrow$ $2\left(3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}\right).$
In this case,
$3\xi^{2}\kappa_{{\rm tot}}^{2}-8\xi\langle\hat{P}^{4}\rangle_{12}+\langle\psi_{1}|\psi_{2}\rangle\langle\hat{P}^{4}\rangle_{12}\rightarrow 3\langle\hat{P}{}^{2}\rangle^{2}-3\langle\hat{P}^{4}\rangle=3{\rm Var}(\hat{P}^{2}).$ (143)
${\rm Var}(\hat{P}^{2})$ is generally nonzero unless the point-spread function $\psi(x)$ is approximately an eigenstate of $\hat{P}$ or equivalently an uniformly-distributed point-spread function over the position space. But this is determined by the property of the point-spread function of either optical source, not by the difference between the two point-spread functions, so there is no new critical quantity determined by the difference of the two point-spread functions from the fourth-order terms in $\mathcal{H}_{d}$.
In summary, we can identify $1-\langle\psi_{1}|\psi_{2}\rangle$, $\chi$ and $\frac{1}{2}-\xi$ as the critical quantities that are determined by the difference between the point-spread functions and will compete with each other and with the ratio $\epsilon$ between the difference and the total of the photon number in determining the validity and the estimation precision of superresolution.
E.2 Relation between critical quantities
When considering the competition between the critical quantities obtained above in determining the validity of superresolution, it should be noted that those quantities are not independent of each other but closely related, and we need to find out the relation between those quantities which will impose constraints on the range of estimation precision for the separation.
Assume that the difference between $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ is small. To facilitate the following computation, we first construct an orthonormal basis for the subspace spanned by $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$,
$|\psi_{1}\rangle,\;|\psi_{1}^{\bot}\rangle=\frac{|\psi_{2}\rangle-\delta|\psi_{1}\rangle}{\sqrt{1-\delta^{2}}}.$ (144)
$|\psi_{2}\rangle$ can be decomposed onto the basis states $|\psi_{1}\rangle$ and $|\psi_{1}^{\bot}\rangle$ as
$|\psi_{2}\rangle=\alpha|\psi_{1}\rangle+\beta|\psi_{1}^{\bot}\rangle.$ (145)
It is straightforward to verify that
$\alpha=1-u,\;\beta=\sqrt{2u-u^{2}},$ (146)
where
$u=1-\langle\psi_{1}|\psi_{2}\rangle.$ (147)
Now, substituting $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ (145) into $\langle\hat{P}{}^{2}\rangle_{2}-\langle\hat{P}{}^{2}\rangle_{1}$, one can have
$\chi=$ $\frac{\langle\hat{P}{}^{2}\rangle_{2}-\langle\hat{P}{}^{2}\rangle_{1}}{\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}},$ (148)
$=$ $\frac{(u^{2}-2u)\langle\psi_{1}|\hat{P}^{2}|\psi_{1}\rangle+2(1-u)\sqrt{2u-u^{2}}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle+(2u-u^{2})\langle\psi_{1}^{\bot}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle}{(u^{2}-2u+2)\langle\psi_{1}|\hat{P}^{2}|\psi_{1}\rangle+2(1-u)\sqrt{2u-u^{2}}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle+(2u-u^{2})\langle\psi_{1}^{\bot}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle}.$
In the above equation, $u$ is a small parameter when the two point-spread functions are close to each other. The numerator ranges between $O(u^{\frac{1}{2}})$ and $O(u)$, and the denominator ranges between $O(1)$ and $O(u)$, so the order of $\chi$ is between $O(u^{1/2})$ and $O(u)$.Similarly, for $\frac{1}{2}-\xi$, we note that
$\langle\hat{P}{}^{2}\rangle_{12}$ $=\langle\psi_{1}|\hat{P}^{2}|\psi_{2}\rangle,$
$=(1-u)\langle\psi_{1}|\hat{P}^{2}|\psi_{1}\rangle+\sqrt{2u-u^{2}}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle,$ (149)
so,
$\frac{1}{2}-\xi=$ $\frac{\frac{1}{2}\left(\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}\right)-\langle\hat{P}{}^{2}\rangle_{12}}{\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}},$ (150)
$=$ $\frac{\frac{1}{2}u^{2}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}\rangle-u\sqrt{2u-u^{2}}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle+\frac{1}{2}(2u-u^{2})\langle\psi_{1}^{\bot}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle}{(u^{2}-2u+2)\langle\psi_{1}|\hat{P}^{2}|\psi_{1}\rangle+2(1-u)\sqrt{2u-u^{2}}\langle\psi_{1}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle+(2u-u^{2})\langle\psi_{1}^{\bot}|\hat{P}^{2}|\psi_{1}^{\bot}\rangle}.$
The numerator ranges between $O(u)$ and $O(u^{2})$, and the denominator ranges between $O(1)$ and $O(u)$, implying that $\frac{1}{2}-\xi$ scales between $O(u)$ and $O(u^{2})$ when the two point-spread functions are close to each other.
In order to make the critical quantities comparable, we denote
$u=1-\langle\psi_{1}|\psi_{2}\rangle=b\widetilde{d}^{h},\;\frac{1}{2}-\xi=a\widetilde{d}^{f},\;\chi=c\widetilde{d}^{e}.$ (151)
According to the above ranges of $\chi$ and $\frac{1}{2}-\xi$ in terms of $u$, we can immediately obtain the constraint condition on the exponents $h$, $f$ and $e$,
$h/2\leq e\leq h\leq f\leq 2h.$ (152)